### v2

# Attention is All They Need: Exploring the Media Archaeology of the Computer Vision Research Paper

Abstract.
Research papers, in addition to textual documents, are a designed interface through which researchers communicate. Recently, rapid growth has transformed that interface in many fields of computing. In this work, we examine the effects of this growth from a media archaeology perspective, through the changes to figures and tables in research papers. Specifically, we study these changes in computer vision over the past decade, as the deep learning revolution has driven unprecedented growth in the discipline. We ground our investigation through interviews with veteran researchers spanning computer vision, graphics and visualization. Our analysis focuses on the research attention economy: how research paper elements contribute towards advertising, measuring and disseminating an increasingly commodified “contribution.” Through this work, we seek to motivate future discussion surrounding the design of both the research paper itself as well as the larger sociotechnical research publishing system, including tools for finding, reading and writing research papers.
Several teaser images from computer vision research papers, including elements like various 3D models of human faces and bodies, images with bounding boxes and distorted images.
1. Introduction
The research paper has many meanings. It is a contribution to knowledge, often responding—according to formal and informal rules of scientific discourse—to a decades-long conversation. It signals to peers, funding agencies, future employers, and future students that the authors are knowledgeable about a topic and actively studying it. It is a unit of productivity which indicates progress towards a PhD, tenure, or promotion. But beneath all of those layers of interaction, it is a digital media artifact, an element of the sociotechnical research publishing system, with a visual form that both influences, and is influenced by, its discipline.
We propose thinking about the research paper not as a neutral medium for disseminating textual scientific content, but as a designed media artifact with which researchers—its “users”—interact with digitally. For example, we might glance at its title in a list of papers, click on it, get hooked by its first page figure, scroll through its pages in a PDF viewer, scanning a table of results. We may also print the paper and carry it with us, write in its margin and leave it on a desk, and then find it months or years later. In all of these interactions, the design of both the visual and textual content, as well as the design of the larger technical system, which produced the PDF, delivered it to us, rendered its image to a screen and transferred it to paper, affects and shapes our interactive attentive experience. Moreover, research papers are cultural objects shaped by disciplinary culture. Bruno Latour, for example, writes extensively about the social nature of the research paper and the important rhetorical role its figures and tables play in creating scientific truth (Latour, 1987). Taking this further, we view the research paper as a media object (Manovich, 2002), an assemblage of designed technologies—like rendered data visualizations, LaTeX, and the PDF file format—viewed on a screen or printed on paper, which have technological affordances and, as we will discuss, accessibility concerns.
In this paper, we study the changes to computing papers in one area—computer vision—over the past decade. While changes occur over time in the research papers of every discipline, computer vision is a particularly information-rich site (Patton, 1990, p.242) for understanding the paper as a media artifact both because of its inherently visual nature and its unprecedented recent growth. Over the past decade, the “deep learning revolution” has transformed several fields of computer science altering both the fields themselves and the way researchers and practitioners feel about them (Su and Crandall, 2021). In computer vision specifically, Su and Crandall find “a general mood of malaise” (Su and Crandall, 2021) which permeates the field. We argue that this malaise is a symptom of increased commodification of scholarly attention, which is materially documented in the changing design of research papers.
By attention, we evoke the attention economy, first proposed by Herbert Simon (Simon, 1971) and applied to the internet by authors such as Goldhaber (Goldhaber, 1997) and Davenport and Beck (Davenport and Beck, 2001). An attention economy in the workplace has been discussed in CSCW, for example in the work of Yardi et al. on the attention economy of internal corporate blogs (Yardi et al., 2009) or Kraut et al. on the attention economy of email marketing (Kraut et al., 2002). To understand its commodification, we turn to the post-Marxist analysis of Claudio Celis Bueno (Bueno, 2017). Bueno frames online attention, not just as a scarce resource, but as a form of labor—human activity which has the potential to generate surplus value. We claim that this theory of attention as labor applies to the attention paid to academic research as well, and helps to explain recent changes to computer vision.
Research papers in computing have had a rich history of debate (e.g., conferences vs. journals (Vardi, 2009)). In the field of HCI and its subfields, including CSCW, debates continue to arise regarding the form and content of the research paper (Geiger, 2019; Wallace et al., 2017). Canonical works like Dourish’s “Implications for Design” comment on the relationship between scholarly publishing and larger issues of power in disciplinary culture (Dourish, 2006). Sometimes, new formats arise in response; for instance, the pictorial format addresses the need to foreground visual imagery as a significant contribution to HCI (Blevis et al., 2015; Blevis, 2016). In HCI, these debates exist both as meta-scholarship, shaping the practices of a discipline, and as subject matter, studying the research publishing system as a technological platform.
Perhaps the biggest hint that the design of computer vision papers has changed with the rapid changes in its disciplinary culture is scholarly humor. In the 2010 computer vision satire paper “Paper Gestalt” (Von Bearnensquash, 2010), the pseudonymous authors “take the simple intuition that the quality of a paper can be estimated by merely glancing through the general layout, and use this intuition to build a system that employs basic computer vision techniques to predict if the paper should be accepted or rejected,” (Von Bearnensquash, 2010) and suggest that this system might replace the peer review process. The visual design of the paper parodies the qualities it observes, including unnecessary complex equations and long algorithms.
But in 2018, Jia Bin Huang published a sequel, “Deep Paper Gestalt” (Huang, 2018), which updates both the methods and the jokes. Juxtaposing these two papers gives a glimpse of how the style of computer vision research papers has changed in just eight years. Instead of using overly complex algorithms and equations, Huang proposes a benchmark dataset, CVPG (the “Computer Vision Paper Gestalt” benchmark), and uses deep learning to significantly outperform the “hand-crafted features” of “Paper Gestalt” (Von Bearnensquash, 2010). Then, it presents an unnecessarily dense table and gratuitously large figure of class activation heatmaps to show the page regions which predict good papers. Though satire, these papers raise real questions about the research paper as a media artifact.
In the following sections, we examine historical computer vision research papers from a media archaeology perspective, focused on the way their visual style has developed over time. To better understand these developments in context, we also report results of interviews with twelve veteran researchers whose work spans computer vision, graphics, and visualization. Our approach focuses on elements of the research paper and larger publication system based on their roles in advertising, measuring, and disseminating an increasingly commodified “contribution.” We find that as academic discourse has moved online, the limiting factor on publishing has shifted from printing costs to the attention of other researchers, which has changed both the culture of computer vision, as well as the design of its research papers. While this finding echoes existing discourse surrounding “publish or perish” (Garfield, 1996; Lee, 2014), the “marketization” of academic discourse (Fairclough, 1993), use of “hype” in science communication (Soto-Sanfiel et al., 2023) or promotional qualities in titles and abstracts (Hyland, 2023), we emphasize the novelty of our approach. To the best of our knowledge ours is the first connection between these discussions of scholarly discourse and the designed visual artifact of the research paper.
2. Related Work
Our inquiry sits at the intersection of several scholarly conversations in and around HCI and CSCW, including data mining of research paper figures, studies of digital media objects, designing for readers and writers of research papers, and studies of the culture of computer vision.
2.1. Research papers as media objects
The fascinating satire paper “Paper Gestalt” (Von Bearnensquash, 2010) is both a related work and a primary source in our inquiry. It claims, facetiously, that “the quality of a paper can be estimated by merely glancing through the general layout” and uses machine learning to predict whether papers will be accepted or rejected from their page images. The (unknown) authors identify that certain features, like sophisticated math and large figures, are predictive of acceptance while large, dense tables are predictive of rejection. A sequel, Jia-Bin Huang’s “Deep Paper Gestalt” (Huang, 2018), updates this methodology using deep learning. These papers are a commentary on the the reality of the modern computer vision community: a reliance on visual performance to ensure successful dissemination and influence of papers.
There is also work surrounding the analysis of research papers in document image analysis (Davila et al., 2020) and data visualization (Wu et al., 2021; Chen et al., 2021). These methods, however, treat the figure as a neutral data source, where there is an objective correspondence between the original data and the figure for each chart type. More recent work has started to, for example, see data visualization as not merely objective reporting of facts but as a communicative medium that has affective impact (Lee-Robbins and Adar, 2022; van Koningsbruggen and Hornecker, 2021). We respond to this literature by pointing out the cultural layers which may interfere with the neutrality and machine-readability of visualization across disciplines.
2.2. Studies of digital media objects
There has been growing interest in analyzing code and its visual results as media objects: artifacts which trace the development of a media culture over time, like wax cylinders or old television sets. Jacob Gaboury’s Image Objects: An Archaeology of Computer Graphics (Gaboury, 2021) takes a media archaeology approach to five such objects in the history of computer graphics: the hidden surface problem, the frame buffer, the virtual teapot, object orientation, and the GPU. In 10 PRINT CHR (Montfort et al., 2014), Montfort et al. study a line of BASIC code used to generate a random maze and use it as an entry-point to explore the cultural history of the Commodore 64 as well as topics including mazes, grids and randomness. These works live within the space charted by Lev Manovich’s now-canonical The Language of New Media (Manovich, 2002), which frames “new media” as visual artifacts produced by textual code, and argues that new media demands new theoretical perspectives.
2.3. The research paper as a site for interaction design research
Closer to CSCW, there are several studies of the research paper as a site for social computing research and designs for digital literature reviewing: Chang et al. introduce CiteSee, a tool for visually augmenting citations in PDF documents based on the reader’s prior research activity (Chang et al., 2023). Qian et al. describe commercial tools for PDF management as “iTunes for papers,” as they treat the research paper as a fundamental unit of scholarship, and propose a literature review tool to break down papers further into collections of specific grounded claims (Qian et al., 2019).
This research community is also interested in designing for the writing process: Head et al. study ways of augmenting digital documents with definitions of terms and symbols to improve readability (Head et al., 2021), and how authors improve readability by augmenting the visual design of their equations (Head et al., 2022). Manzoor et al. develop a LaTeX editor extension to improve the accessibility of LaTeX for authors with visual impairments (Manzoor et al., 2018) and Hara et al. develop a system for generating Braille documents from mathematical expressions written in LaTeX (Hara et al., 2000). Gobert and Beaudouin-Lafon conduct a study of LaTeX users and design an extension for VSCode that uses transitional representations of document objects like tables to improve the editing experience (Gobert and Beaudouin-Lafon, 2022). Haber et al. study how groups of coworkers interact differently when using physical vs. virtual documents (Haber et al., 2014). While we do not design or develop any technical tools, our work provides additional evidence for the importance of further design studies of research paper reading and writing.
2.4. The culture of computer vision
Motivated by the dangers of algorithmic discrimination and safety concerns in systems relying on computer vision algorithms, several recent papers have studied the culture of the computer vision research community and its understanding of data and truth, often using research papers as texts for analysis. Su and Crandall study the emotional state of the computer vision community, finding that the deep learning revolution and subsequent growth has had a profound effect on its culture, and leading both to excitement regarding progress as well as to isolation and malaise (Su and Crandall, 2021). Denton et al. use a discourse analysis approach to study the history of the ImageNet dataset through the research papers and presentation slides of Fei-Fei Li (Denton et al., 2021). Scheuerman, Denton, and Hanna study a corpus of 500 papers describing computer vision datasets and analyze the values implicit in their writing: efficiency, universality, impartiality, and model development over dataset development (Scheuerman et al., 2021). While our work also analyzes computer vision though its research papers, our interest is in studying these papers as primarily visual media.
A wide variety of authors have written critically about the culture of data collection and dataset use in machine learning. For example, Sculley et al. and Ethayarajh and Jurafsky critique the concept of benchmarks and leaderboards in machine learning, arguing in different ways that steadily increasing scores do not always correspond to progress (Sculley et al., 2018; Ethayarajh and Jurafsky, 2020). For a more comprehensive survey please see Paullada et al. (Paullada et al., 2021). Our work here, which does not investigate data collection or dataset usage directly, nevertheless echoes these themes.
3. Methods
Our inquiry began through visual analysis of historical computer vision research papers. We found the changes to the visual design of these papers surprising, particularly the increasing prevalence of highly complex figures, and decided to investigate further by conducting both semi-structured interviews with researchers who had been active in computer vision, graphics, and visualization for several decades, and computational analysis of a corpus of research papers published between 2013 and 2021 in the IEEE Conference on Computer Vision and Pattern Recognition—the highest ranked publication venue in computer science and third highest in all of science, according to Google Scholar (goo, 2024). We study this conference, rather than similar conferences in machine learning or natural language processing, because of the importance of visual presentation in computer vision. Along the way, our visual analysis, interviews and computational findings condensed into a media archaeology approach.
3.1. Media Archaeology
Inspired by recent work on the history of computer graphics (Gaboury, 2021), we examine historical research papers through the lens of media archaeology. By “archaeology,” this method refers to Foucault’s concept of archaeology as the “uncovering of the archive” (Foucault, 1969, p.131), rather than physical excavation. Media archaeology, as defined by Huhtamo and Parikka, “rummages textual, visual and auditory archives as well as collections of artifacts, emphasizing both the discursive and material manifestations of culture” (Huhtamo and Parikka, 2011, p. 3). In other words, we study digital media objects as components of a media culture; a system of practices and meanings which structure our interpretation (Rose, 2016). We also consider their digital materiality as physical patterns of bits on hard drives, produced through the labor of researchers. By drawing attention to material culture, media archaeology is in dialogue with theories of historical materialism, which consider economic factors, production and consumption of goods and services, as the driving force of history, rather than revolutionary ideas or progress narratives (Gaboury, 2021, p. 5).
To illustrate these concepts, consider Figure 2(d), a teaser image from a paper in the proceedings of the IEEE Winter Conference on Applications of Computer Vision (WACV) (Kong et al., 2022). This figure is split into 5 similar subfigures, with different combinations of black lines and teal dots. With the requisite disciplinary knowledge, a viewer can interpret the meaning of the figure: the five subfigures represent an occluded image and four attempts to reconstruct the wireframe of the original using different automated methods. Using our understanding of the task, we clearly see that the fifth subfigure labeled “ours” is the best, because its dots and lines align with the geometry of the depicted room. These inferential steps require a degree of initiation in computer vision, an understanding of both the goal of a research paper and the system of meaning that these papers use. The goal of media archaeology is to dissect these inferences and identify how their layers of meaning developed over time.
In addition to our computational analysis and detailed study of the CVPR proceedings, we “rummaged” research papers from the IEEE, Computer Vision Foundation, and ACM digital archives, including the IEEE International Conference on Computer Vision (ICCV), European Conference on Computer Vision (ECCV), IEEE Transactions on Pattern Analysis and Machine Intelligence (PAMI), and the ACM Speical Interest Group on Graphics and Interactive Techniques conference (SIGGRAPH). We examined ICCV, ECCV, and PAMI in addition to CVPR because they are the four most influential computer vision venues according to Google Scholar,111https://web.archive.org/web/20220709182958/https://scholar.google.com/citations?view_op=top_venues&hl=en&vq=eng_computervisionpatternrecognition and SIGGRAPH due to its historical significance and prevalence in our interviews. We turned to these venues to build context for the results we were seeing in CVPR. While we did not systematically code the enormous number of papers published in these venues, we collected dozens of screenshots of interesting figures and tables and discussed their visual style in weekly meetings, integrating interview and computational results which eventually coalesced into historical narratives.
3.2. Interviews
To augment our visual analysis, we conducted interviews with veteran computer vision researchers. As there are a relatively small number of eligible participants, we recruited individuals via email and in-person at the CVPR 2022 conference. Participants are listed by years of industry and academic experience in Table 1. This sample is highly non-representative, showing both survival bias based on who remains in the field over a long period of time and selection bias based on who agreed to an interview, so it does not serve as the source of our claims. Instead, it provides important social context to our visual analysis.
We conducted interviews in person and over the Zoom videoconferencing platform between March and August 2022. To avoid participants historicizing or theorizing themselves, we asked each participant to discuss a specific research paper from their early career, and asked them to explain the different elements and tell us stories about the writing process for that paper. Additionally, we asked each participant how they read research papers at that time of that paper, and how those reading and writing processes differ from those of the present. This method of interviewing mirrors that of previous studies examining media artifacts such as websites (Chen et al., 2017; Goree et al., 2021).
After transcribing the interviews, we engaged in a process of iterative memoing, integrated into our weekly meetings where we compared examples from historical research papers and interview excerpts which shared common themes. We focused our analysis on identifying patterns from the context that participants provided while telling stories about their writing, as well how participants thought about their papers in relation to changing disciplinary practices, rather than simply analyzing the papers participants discussed, or discussing our participants’ analysis of their own work.
| Participant | Discipline | Years Industry | Years Academia |
| P1 | CV,G | 9 | 33 |
| P2 | G,V | 3 | 17 |
| P3 | CV | 0 | 42 |
| P4 | CV | 21 | 3 |
| P5 | CV | 15 | 11 |
| P6 | CV | 0 | 10 |
| P7 | CV | 0 | 36 |
| P8 | CV | 0 | 16 |
| P9 | CV | 0 | 19 |
| P10 | CV,V | 0 | 22 |
| P11 | V | 11 | 0 |
| P12 | CV,G | 0 | 23 |
10 participants come from computer vision, 3 from graphics and 2 from visualization. 11 have academic experience, ranging from 10 to 42 years, 5 have industry experience ranging from 3 to 21 years. Facetiously, the highest number in each column is bolded.
3.3. Computational Analysis
As both qualitative approaches rely on the analysis of specific examples, we used supplementary quantitative analysis to verify that phenomena we observed in the interviews and media archaeology were actually as pervasive as participants seemed to think. Specifically, we were interested in whether teaser images, figures and tables were becoming more prevalent in CVPR papers over time, and whether more titles were following a particular format containing an acronym followed by a colon. To answer these questions, we collected open access PDF files from thecvf.com published in CVPR 2013 to 2021 and article metadata from ieeexplore.ieee.org for CVPR 1992 to 2020. We manually inspected PDFs from thecvf.com to count teaser images, then used the Linux pdftotext tool combined with regular expressions to count figures and tables. From the IEEExplore metadata, we used regular expressions to parse paper titles, treating a word with more than two capital letters as an acronym, and an acronym as unique if it only appears once in a given year of data.
4. Findings
Six subfigures. (a) shows Randy Pausch’s 1996 teaser image, a large color figure of an Aladdin-themed virtual reality ride. (b) shows a 2021 visualization paper from TVCG which uses superhero imagery to illustrate its design approach. (c) shows the ACM conference template, with its stock baseball teaser image. (d) shows a paper from WACV 2022 with a teaser showing an image of a room containing a missing image region and four attempts to locate the edges of the room. The last method labeled “ours” clearly has the best results. (e) shows an input image of a small figurine, possibly Moses, and five competing methods to predict saliency maps over the image. Again, the last method, marked “ours” does best. (f) a still from a paper towel TV advertisement showing “Bounty select-a-size” effectively cleaning up a spill while “Next leading 2 ply” fails. The image also reads “Laboratory demonstration. 3 passes, 3.5 sec elapsed time, 1 stp caramel, 0.5 psi.”
Through our analysis, we found that several elements of the contemporary computer vision research paper serve as material traces of the disciplinary change which took place over the course of the 2010s. We find that papers are written to promote a specific contribution which is increasingly commodified. By “contribution,” we mean the model, algorithm, method, dataset or other system that the paper offers to readers. By commodified, we mean reduced to its exchange value in terms of the attention it generates. As Simon theorized (Simon, 1971), and authors including Goldhaber (Goldhaber, 1997) and Davenport and Beck (Davenport and Beck, 2001) have applied to the web, when “capital, labor, information and knowledge are all in plentiful supply” the limiting factor is the attention of consumers, which begins to be treated like a commodity (Davenport and Beck, 2001).
However, this concept of the attention economy presupposes that attention is valuable, which is unhelpful for understanding why academics seek the attention of their peers instead of the larger research community or general public. To better understand why academic attention is uniquely valuable, we employ Bueno’s explanation, that online attention, in addition to a commodity, is a form of labor (Bueno, 2017). This claim relies on Romano Alquati’s concept of “valorization information” via Pasquinelli (Pasquinelli, 2015). Alquati, in his study of an Olivetti computer factory, found that workers contribute to mostly-automated production processes through their micro-decisions, which served as a feedback-control signal for the production process. Bueno argues that attention can provide a similar signal. Specifically, by monitoring attention, companies can forecast consumer behavior and more efficiently follow market trends. By doing so, these companies extract surplus value from online behavior data (Bueno, 2017, Ch. 2). We believe a similar process may be at play in academic publishing, where authors can leverage highly specialized early attention to increase the prestige value of their research. This process can take the form of feedback, where authors improve their work based on an early online response, or simply where early viewers find a paper interesting and share it, judging it as valuable, putting it in front of more eyes and increasing the chance that it will reach the right audience. Eventually, that attention translates into citations, jobs, students, grants and other benefits of academic prestige.
We study three aspects of contemporary research papers: the teaser image, the results table and the high resolution figure. We have grouped our results into thematic sections surrounding these three elements in relation to the concept of a paper’s contribution.
4.1. Advertising the Contribution
4.1.1. Teaser images
The first element of research paper design that we examine is the “teaser image,” a large first-page figure which summarizes the paper. Several examples of teaser images are shown in Figures 1 and 2. These figures are functionally similar to the trend of “visual abstracts” (Ramos and Concepcion, 2020) and “table of contents images” (Buriak, 2011) in the biomedical sciences, however unlike these forms which are graphical summaries of the text, teaser images are part of the main paper and are usually a visualized system output that the authors want to show. Teaser images have been steadily gaining popularity within CVPR, as shown in Figure 3 left. For this analysis, we defined the teaser image as a first-page figure which covers the entire width of the page, not just a column.
Two subfigures. Left shows a graph with “fraction of papers with teaser figures” on the Y axis and “Year” on the X axis. The line increases from 2% in 2013 to 12 % in 2021. Right shows a graph with four lines, representing four quantities which increase between 1992 and 2020: “Title: with colon” increases from 10% to 35%, “UA (unique acronym)” increases from 7% to 17%, “AFbC: Acronym Followed by Colon” increases from nearly 0% to 22% and “UAFbC: Unique Acronym Followed by Colon” increases from nearly 0% to 15%.
P12 ascribes the teaser image to the famous graphics researcher Randy Pausch: “Randy Pausch…takes credit for putting the teaser image in SIGGRAPH papers…he claimed that he did a paper where half the first page was a teaser image and after that that became the norm where people started always putting these images.” The paper in question is pictured in Figure 2 (a). While Pausch did not invent the large first-page figure (Tim Berners-Lee famously used one in his 1989 proposal for the world wide web, for example (Berners-Lee, 1989)), he did publish the first paper in SIGGRAPH with this layout, and it quickly became popular. P2 and P10 point to the quick adoption of teaser images, which even became institutionalized in the SIGGRAPH template, which is now the template for all ACM conferences (Figure 2 (c).
P12 defines the teaser as a visualization of either a result or a system and explains that it has spread to many other conferences because of the way it attracts attention: “It’s a trailer. It’s to get people in…I think it’s a very compelling way to convey what the paper does.” For her, the teaser is a highly effective innovation which improves research papers. P2 echoes that sentiment: “They just made the papers look good! I mean, It’s much more memorable and there are some papers still today that I don’t remember the title, but you see the picture and you’re like, oh yeah, that’s the Randy Pausch paper on the VR for whatever, right? So yeah there’s a few of those that are just like, really iconic first pagers.” P2 is referencing the same Pausch paper as P12 (Pausch et al., 1996), highlighting its memorability.
The teaser image is a trailer, a hook; it advertises the paper to the potential reader. The authors want to promote their paper and showcase the best results they can because the sheer number of and easy access to papers has made it harder to stand out. The visual organization of these figures echoes that theme. Notice the commonalities between the two teaser images in Figure 2 (d) and (e). Both show the output of several algorithms attempting to solve the same problem, but they depict the results in such a way that their method is clearly best. This visual effect mirrors that of advertisements, which depict two competing products in action (Figure 2 (f)). While the experiments in research papers are far more sophisticated than those in advertisements, they share a common visual design. Along these lines, we find a recurring theme of teaser images which rely on the iconography of the 3D-reconstructed, brightly colored human body as a particularly compelling visual. Several examples are shown in Figure 1. Human faces and bodies have been found to attract consumer attention to advertisements (Wilkinson and Light, 2011; Ohme et al., 2011); a similar principle may be motivating computer vision authors.
We observe a similar advertising quality in titles. Figure 3 right shows the rise in popularity of a particular title construction where an acronym, which is usually the name of a model, is followed by a colon: “HOPE-Net: A Graph-Based Model for Hand-Object Pose Estimation” (Doosti et al., 2020) or “DeMoN: Depth and Motion Network for Learning Monocular Stereo” (Ummenhofer et al., 2017). These names both signal that the paper proposes a new model and brands the paper with a short, memorable name which is often a cultural reference or clever pun. An especially cheeky example is Joseph Redmon and Ali Farhadi’s 2017 paper “YOLO9000: Better Faster Stronger” (Redmon and Farhadi, 2017), which improves upon an earlier model called YOLO (short for You Only Look Once) (Redmon et al., 2016), and references a variety of memes from the mid-2010s.222YOLO itself was a meme, short for “you only live once,” the number 9000 references a famous quote from the television show Dragon Ball Z and the phrase “better, faster, stronger” references a Daft Punk song. While unusual in computer vision, this kind of nerdy referential humor has been observed in other areas, for example, web design (Goree et al., 2022).
4.1.2. Beyond the paper: arXiv, social media and videos
Why would many authors begin to engage in these teaser figure and title practices over the course of the 2010s? We find that it has resulted from a change in medium through which research reaches its audience, from conference and journal proceedings onto social media and preprint websites. P4 describes how his students have strategies to get their papers noticed on the preprint site arXiv: “if they take your submission on Thursday and it goes up on the arXiv most recent publications list it’s up for all of Friday, Saturday, and Sunday…so it gets more days of exposure…It improves your odds that somebody will notice your thing.” P8 describes the pressure on her students to promote their work: “social media, like promoting one’s research has become such a big thing and I think students…are realizing, oh it’s not enough for me to, y’know, come up with a paper, post it on arXiv, get it accepted. I also need to tweet about it. It’s frankly quite exhausting…[before] the only way you promote your paper is it shows up at the conference and hopefully, y’know, some famous computer vision researchers will come up and look.” She talks about how her advisor would bring his friends over to her poster, and that was all the advertising she needed to get noticed, but that strategy no longer works as well in the modern crowded field of computer vision.
Again this pattern resembles one from computer graphics. P12 describes the importance of videos in SIGGRAPH: “in graphics the conventional wisdom has been it’s useful to have, not only do the paper…but to also show a video that really highlights and explains the work. And if you do a good job of explaining it, people find it compelling.” These videos were considered part of the paper submission and peer reviewed alongside the text, which led authors to invest heavily in the quality of their videos.
That attention translates into citations, as well as other benefits. P11, who works in an industry lab that hires recent PhDs, explains why his group publishes at all: “exposure is actually really important. If you want to attract really top level talent, then having zero published papers is really going to work against you, right? Particularly if you’re looking at people who are in positions that are for doing some sort of active innovation.” For this group, the publication is primarily a way to gather attention, they derive value from the attention paid to their work through hiring. Second, attention helps researchers to identify the value of their own work. P7 talks at length about his desire for impact. He switched fields from computational complexity to computer vision because there were more active researchers, and he could not tell if his work was having an impact in a small field. P9 tells the story of how attention from a senior scholar improved his relationship with his advisor: “that email from [scholar], that’s the beginning of our of my great relationship with my advisor was that one little email, because within five minutes [advisor] was up in the office, ‘Oh, you got an email from [scholar]? Very good. Very good. How about a coffee?’ ” Notice that the career benefits resulted, not from the attention itself, but from the record of that attention in the form of an email from a senior scholar.
While the relationship between attention and prestige is not new, the unprecedented growth of computer vision over the past decade places pressure on the attention of senior researchers. There have been ongoing discussions in the computer vision community on how to react to this growth. Some of the proposed ideas are recorded in the passed motions of the annual meeting of the IEEE pattern analysis and machine intelligence technical committee (PAMI-TC), the governing body of the computer vision research community. In 2021-2022, motions changed the rules regarding double submission to reduce burden on reviewers, introduced punitive measures for missing or poor quality peer reviews from authors and banned discussion of submitted papers on social media to avoid influencing peer reviewers, the last of which was repealed in 2023 (pam, 2024; Black, 2021). The rationale statement for the social media ban specifically highlights attention: “Groups with large followings and the resources to mount visible social media promotions received significant attention for work that is under review. Reviewers are exposed to this work and the attention it receives can bias their judgment—if so many people on social media are excited, mustn’t it be good?” (Black, 2021) In addition to highlighting the valorizing role of attention, this statement showcases the view of peer review as an unbiased process for evaluating the quality of research.
We see two additional key themes starting to emerge. First, there are parallels between computer vision in the 2010s and computer graphics in the 1980s and 1990s, which many of our participants pointed out (P1, P2, P6, P9, P10, P12). Both disciplines rapidly grew due to industry investment, from the tech industry in vision and the entertainment industry in graphics, which created an attention economy, forcing papers to go above and beyond to attract attention. Second, research work is being commodified—treated as interchangeable, given some measure of its value. In computer vision, that means the particular ideas an author proposes are less important than their ability to grab attention and advance the author’s career.
4.2. Measuring the Contribution
Two subfigures, both line plots. Left shows the percentage of papers with a figure remaining constant around 99% while the percentage of papers with a table increases from 77% in 2013 to 97% in 2021. Right shows the average number of figures and tables per paper. Figures are constant between 6 and 7 while tables increase from 2 in 2013 to 4 in 2021.
Today, the “table of results” in a computer vision paper fulfills a central function in both the written argument and the peer review process: evaluation. When proposing a new method for solving a vision problem, the authors must demonstrate that it works at least as well, if not better than, “SOTA” (state-of-the-art) existing methods. These tables often contain the values of standard evaluation metrics computed on a benchmark dataset, like top-1 or top-5 accuracy on the ImageNet test set (Deng et al., 2009) for image classification or mean average precision on the MS-COCO test set (Lin et al., 2014) for object detection.
While it is tempting to treat these tables as textual content, their effect on the paper, we argue, is primarily visual. A key feature of these tables is that they put the best result, which is almost always from the author’s proposed method, in bold. This design feature is essential for readability, as a large table full of numbers is very difficult to interpret. These tables will sometimes also use arrows to indicate whether a column displays a metric where higher or lower numbers indicate better performance. More recently, as these tables have become more complex, authors have developed other readability innovations, like using colored numbers and subscripted or parenthetical percent improvements (Figure 5 (e)). These innovations serve to further visualize the quantified value of the paper’s contribution.
But all of this was not the case a few decades ago. The vast majority of computer vision papers from the 1980s and 1990s rely on mathematical arguments based on pinhole camera geometry and do not contain any quantitative results. Empirical evaluation, if included at all, was primarily qualitative, in the form of figures showing sample results. As P8 explains, she had a combination of quantitative and qualitative evaluation in her paper from 2003, which was unusual: “quantitative evaluation, you know, back in 2003 was still kind of in its infancy…I’m not sure that this [2003] paper has basically any comparison to competing methods which probably would be required today.” P3 explains that he was primarily concerned in 1990 with showing test examples to demonstrate his algorithm’s effective handling of edge cases. P9 explains that in 1999, showing example output of his system was sufficient: “instead of [Amazon] Mechanical Turk you just have the reviewers just eyeball the images.” In the satirical 2010 “Paper Gestalt” (Von Bearnensquash, 2010) paper, which attempts to use computer vision methods to distinguish between good and bad papers, large confusing tables were identified as a key feature of bad papers, not an essential feature of good ones.
So how did computer vision transform from a mathematical discipline based on geometry to an empirical, quantitative discipline based on benchmarks? This transition was gradual: we can see its seeds as early as a debate at ICCV 1999, referenced by P9, between Jitendra Malik and Olivier Faugeras (Triggs et al., 2000). In that debate, Malik argued that computer vision should focus more on probabilistic modeling and perception, rather than methods based in geometry, while Faugeras rejected empirical computer vision as unfalsifiable, advocating for the mathematical rigor of geometry.
Six subfigures. Each shows a table full of numbers, where some of the numbers are in bold. (a) shows a table split into sections with double lines where one number in each row is bolded. (b) shows a simpler table where all of the numbers in one column are bolded. (c) shows another simple table where one row has an italicized label and a citation, while two of the other rows have numbers in bold. (d) shows a highly complex table with dozens of rows split into several sections with horizontal lines and several numbers in bold. (e) shows a dense table with a grayscale background, alternating in tone to differentiate the lines, and numbers in both blue and red. (f) shows a table with many rows split into sections with double horizontal lines, two numbers in bold and several cells with infinity symbols.
The publication of Krizhevsky, Sutskever and Hinton’s “ImageNet Classification with Deep Convolutional Neural Networks” in 2012 (Krizhevsky et al., 2012) marks a turning point for empirical evaluation. This paper is historically significant for setting off the deep learning revolution, and its design and writing served as a foundation for the thousands of deep learning-based computer vision papers that followed. The paper’s central argument is that several “new and unusual features” lead deep convolutional neural networks to significantly outperform other methods. These features include rectified linear units (ReLU), GPU-based training, and regularization techniques like data augmentation and dropout. Crucially for our story, however, this argument is made by way of a table, shown in Figure 5 (c), with the best performance in bold. Neural network papers are obligated to use empirical evaluation, as there are insufficient theoretical guarantees for these models and they are difficult to evaluate otherwise. Over the following years, many papers followed Krizhevsky’s lead, showing that deep convolutional neural networks outperform existing methods on other central problems like object detection and semantic segmentation. We can see a corresponding increase in both the average number of tables per paper, as well as the fraction of papers containing at least one table in Figure 4. While the prevalence of figures has remained relatively constant, tables have become significantly more common. While only 75% of CVPR 2013 papers had a table, 95% of CVPR 2021 papers did, and the average number of tables per paper has doubled from 2 to 4. While not all of these tables are the kind of results table we’re discussing, the change is striking.
Like teaser images, however, this style of table arises in computer graphics before entering computer vision; see Figure 5 (a) for an early example. Early graphics results tables primarily showed runtime comparisons, rather than accuracy or quality evaluations. These tables are used in computer vision for showing machine learning performance at least as early as 1999 (Figure 5 (b)). Interestingly, tables from this era usually had methods in the columns and different data examples in the rows, in contrast to the later tables which have evaluation metrics in columns and methods in rows. This swap also aided readability, as it is easier to scan vertically than horizontally (Luckiesh and Moss, 1941).
As competition has heated up, results tables have grown. For example, compare the table in Figure 5 (c), from a 2012 paper, to the table in Figure 5 (d), from a 2021 paper. The benchmark remains ImageNet, though performance has surged from 40% top-1 accuracy to over 85%, but the competing state-of-the-art includes dozens of models and differ by only fractions of a percent. Even highly geometric papers, like the example in Figure 5 (f), now involve empirical evaluation.
More broadly, in computer vision there is widespread assumption that research inherently involves competition between technical methods. Several of our participants described academic publication using free market metaphors. P2 explains his research output: “publication was really fast during that time because there was not a lot of competition.” P3 explained that CVPR has become inaccessible to students because “supply and demand” have raised the standards for publication. P4 described vision as “very industrial” and gave examples of techniques that students use to optimize their arXiv submissions to reach as many eyes as possible.
While these tables are buried deep within the paper, it may seem like their effect on attention is minimal. However, a competitive benchmark result attracts attention on its own. Once a researcher has the lead, other papers must compare against their benchmark and cite them. P6 compares benchmarks to arcade game leaderboards: “If it’s an established benchmark, there’s like there’s somebody who has the lead right? Like you would have on like a classic video game arcade machine, right? It’s like I have to just get my initials at the top right? That’s exactly what they’re doing, and that’s frustrating.” P5 and P8 advise their students to avoid working on research which forces them into competition with large companies. P5: “I tell them, don’t work on problems that, you know, a lot of people are working on right now, you can’t possibly compete with Facebook, Google, Amazon because you’re not as computing heavy.” While students could still develop innovative ideas for those problems, they likely would not beat the benchmarks that large companies have set without leveraging similar computing power and would not attract attention.
In summary, the development of the results table shows the transition of computer vision from an applied mathematical discipline to a quantitative-empirical one. As vision started relying on empirical evaluation, it adopted a style of results table which was used for runtime benchmarking in computer graphics, and as evaluation benchmarks became established for vision problems, these tables grew in size and importance. Today, competition on benchmarks is the organizing principle of the discipline, new methods must demonstrate that they are empirically more effective than existing methods to be accepted, and the results table is an essential part of the paper. This element showcases several of the same themes as teaser images: influence from computer graphics, and commodification, in the form of measurable improvements over prior work. It also echoes patterns observed in HCI more broadly regarding quantification: once a phenomenon, in this case the quality of a method, has been measured, that measurement creates and constrains possibilities for action (Pine and Liboiron, 2015).
4.3. Disseminating the Contribution
Six subfigures: (a) a screenshot of a research paper with two scatterplots in the top right corner, occupying a tiny fraction of the page. (b) and (c) are zoomed versions of the scatterplot, revealing that each data point is represented with a digit drawn on the pdf. (d) shows a page with a dense neural network diagram in the top third with tiny mathematical annotations. (e) shows a page with nine ROC curves and 9 pie charts in the top third. (f) shows a page with three figures and a table in half the page area, one figure contains three scatterplots and a line graph, the second contains two matrix heatmaps with tiny labels for each of dozens of columns, and the third contains 25 images of faces.
4.3.1. The PDF and digital proceedings
The final element we explore is the Portable Document Format (PDF) itself, and the way that authors engage with it as a medium. In this section, to avoid confusion, we use “research paper” to refer to printed papers and “research PDF” to refer to the digital version. PDF was a derivative of Adobe’s already-successful Postscript programming language for specifying documents, and a proprietary file format until 2008. In an article from 1998, Kasdorf compares the PDF format to SGML (Standard Generalized Markup Language), and argues that both formats should be used on the web: markup languages for screen-based content and page-based languages for printed content (Kasdorf, 1998).
However, in the context of computer vision research, the PDF is now more of a screen-based format than a printed one. The PDF viewer has steadily replaced paper as the medium for reading research articles, which has led researchers to design their articles for viewing on screens, rather than as physical, printed research papers.
Our participants ascribe this shift to an aesthetic factor: color. P9 explains that a huge draw of a conference like SIGGRAPH was its color proceedings: “SIGGRAPH was always in color. It was absolutely beautiful color. Much more expensive.” P10 points to a turning point where researchers started preparing all of their figures in color: “There was also a turning moment at some point, I think it was around 2004-5-6 something like that. So before that, it was black and white plus color images at the end. Like in a set of separate full color plates as they call them. It was difficult to decide which figures to put in color, since visualizations had to be fully redesigned for black and white.” Once some figures could be in color, authors had to choose which of their figures to leave in black and white, which was a difficult choice. P10 explains his solution: “the only simple solution I found is you don’t decide. You just leave it like this and then they complain, well this is not visible…[so] you insert the URL of your webpage.”
But the existence of URLs, and particularly hyperlinked URLs, points towards another motivating factor for digital proceedings. Before the web, articles were written, submitted, and published on paper. But through the 1990s and 2000s, a series of rapid technological changes took place, transitioning to digital submission and dissemination. The transition to digital proceedings, like other trends, started in SIGGRAPH, which produced the first fully digital proceedings on CD-ROM in 1993 (Brown and Cunningham, 2007). CVPR distributed digital proceedings as early as 2005, allowing color figures without the cost of color printing. P9 explains that in 1999, digital versions of papers were hard to read: “most papers were in postscript and the postscript viewer was kind of sucky. Didn’t really look very good on the screen.” P7 found the transition to PDF freeing: “I’ve seen figures have become a whole lot more common…the PDFs are now in color so you have a whole bunch of other choices.”
One such choice is the high resolution figure, which exploits the scalable nature of vector graphics and the zoom feature of PDF viewers to put more data in a research PDF without exceeding the page limit. Several examples of this trend from CVPR 2021 are shown in Figure 6. While these figures are not overly complex or poorly designed, they are too small to be readable on a 300-dpi printed page.
Several of our participants mentioned these high resolution figures. P6 and P10 are frustrated that they cannot print some PDFs because the figures are too small. P10 ascribes these figures to the increasing complexity of research: “the tendency of grouping more small figures into a kind of collection… I wonder, is this because of the limited page count that people try to avoid the whitespace around the figures and then they put everything in one big figure because then you can save on the whitespace? Or is it simply because really…the techniques that we’re describing nowadays are so, so much more complicated? So you need them to, you know, to, to put the architecture of the whole thing in detail there. Otherwise nobody understands it.” P7, on the other hand, ascribes it to digital submissions: “In ’94, you turned in your paper by actually having a piece of large paper that was probably 24” wide by 30” or something… you had columns of text from LaTeX and then your figures had to be glued in place in what’s called ‘screen,’ right? So you actually have to make sure that they didn’t copy badly.” When papers were submitted by mail, the medium of the “camera-ready” draft was a significant limitation on the kinds of visual content that could be included, but now the only limitation is on the number of pages, so authors use smaller figures to include more content. The small size of these figures, especially in comparison with the large front page teaser image, creates a visual hierarchy, highlighting the differences in purpose between these figures. The teaser attracts attention, the dense figures provide details for those who zoom in, and the impression of details for those who do not.
The digital nature of contemporary research PDFs was a sensitive topic in our interviews. Half of our participants (P1, P2, P3, P4, P5, and P8) expressed attachment to and nostalgia for their printed papers, and were saddened when asked when they stopped reading on paper. For P1, the time spent at the library made the papers physically significant: “[the papers] were all me standing at a Xerox machine with a journal…I didn’t find it as annoying. I found that it gave me something like a kind of a physical connection to the paper. This was sort of like, it’s mine. I’ve watched it go by, you know.” P2 remembers that the SIGGRAPH proceedings “were like flipped to the point where like it’s the threads were like bare.” He’s kept all of the papers he printed during his master’s degree and revisits them from time to time: “I still go back, you know, look at like oh this crap that I thought about these things, right?” Finding and photocopying papers made them precious; he thinks there is “a kind of a correlation between that sense of, like scarcity of it, like how precious the paper was, because it’s so hard to find versus the amount of care you give it.” P4 fondly recalls his advisor taking papers out of a filing cabinet and photocopying them for him: “the amazing thing about your advisor is you have this filing cabinet of stuff that’s already pre-copied, you know, every paper he thought was interesting.”
These sorts of relationships to printed papers are interesting from a design perspective. They echo the findings of Odom et al. that we preserve designed objects which have functional, symbolic and material-aesthetic value (Odom et al., 2009). When a researcher prints out and writes on a paper, the paper gains symbolic significance, a “physical connection” which endures over time. That makes printed pages less disposable than digital ones; there is a more tangible opportunity cost to creating them. From this perspective, digitization is an essential component of commodification, as it separates the research from its paper and thus its role as symbolic object. While reading digitally is significantly more environmentally sustainable than printing, shipping and photocopying paper (Aydemir and Özsoy, 2020), many of our participants who primarily read on screens miss the era of physical papers.
In contrast, P7 prefers PDFs because they allow him to use a screenreader: “I use an app that reads the pdf aloud to me…I can get most of the paper from that, I still have to read equations.” Despite being an advocate for making computer vision papers more accessible for the community, he explains that conferences are reluctant to put accessible equations in their templates because of its impact on file size: “basically the conferences keep using templates that—so there’s a trade off by adding this stuff to your file, your file gets bigger. And it turns out that the package, that one of the packages is pretty good for this, if you do it badly, it blows files up huge…I tried to get the guys from CVPR to use it and it just didn’t happen, right?…I’ve just lived with the fact that it’s not there.”
The juxtaposition between high resolution figures and inaccessible equations is ironic. These figures can have surprisingly large file sizes—the PDF shown in Figure 6 (a) is a stunning 36MB, more than 10 times the size of a typical CVPR PDF. But computer vision proceedings remain inaccessible, justified based on file size concerns. These figures also point to the vestigial nature of the page limit. Originally, page limits were put in place to minimize printing costs. But in the era of online-only proceedings, there is no financial reason to keep papers page-limited. In fact, the main limiting factor on conferences is the number of reviewers, rather than the lengths of accepted papers. As P8 describes: “for the last…five years or more, there has been talk every single cycle, oh, you know, we don’t have enough qualified reviewers. We have way too many submissions. Everyone’s way too overwhelmed to do the reviewing.”
4.3.2. Faster publication through arXiv
The digital materiality of the research PDF affords an alternative publication process that came up in several of our interviews (P4,P5,P8): the preprint server arXiv. P5 says arXiv is a major source of anxiety for his students: “So my students…had the unpleasant experience of, you know, finishing a paper when they’re just about to submit and they saw a paper on arXiv that did almost the same thing. Like, oh my, months of work just went down the drain.” This emotion, the sense of loss when a project becomes unpublishable because someone else got there first, echoes Su and Crandall’s observation of “selective amnesia” (Su and Crandall, 2021), except in addition to old papers quickly becoming obsolete because something else is better, current papers may become obsolete because they are no longer first.
Conferences originally gained popularity in computer vision because they allowed research to make it to print faster than journals. According to P3, “It used to take almost two years from the time a research is done and the PAMI [Pattern Analysis and Machine Intelligence journal] paper would appear. And so the conference has started becoming more and more important at that time. So that’s the origins of ICCV and CVPR conferences becoming far more popular than journals. So during the 1980s and early 90s, the journal was the thing.” But now, if research is primarily disseminated by PDFs posted on arXiv, it seems like the same process is occurring again: conferences behave like journals and arXiv behaves like a conference. As P8 explains, researchers still submit to conferences for prestige: “In order to get promoted in order to get a job, you need that stamp of approval. And it’s a pretty strong signal to get papers accepted to a selective conference…this is really kind of the, you know, biggest marker, you know, the biggest benchmark by which you are evaluated.” As arXiv replaces conference proceedings as the fastest communication medium, the conference is left primarily to evaluate papers.
ArXiv changes the writing process as well. As P9 explains, “I find that the amount of time and effort put into each paper has gone downhill. For my papers, I would put a huge amount of time, especially in the intro…And recently I have been doing less and less of that and basically because the students are saying, ah you know, the new kids they don’t even, they just skip the intro, they just go directly to the method. So I feel like, oh my god, nobody’s even gonna read my beautiful prose!…But also I think because the field progresses faster, papers become obsolete much quicker. So it might be reasonable not to spend so much effort on a single paper if you know that in a year it will be obsolete.” In other words, author writes differently in order to better fit the faster publication system, spending less time on introductions in order to spend less effort on a paper which might quickly become obsolete.
To summarize, the technologies underlying the publishing process in computer vision have changed rapidly over the past three decades. Today, proceedings are published online, and most researchers read research PDFs on a screen, rather than research papers in print. The loss of physical research papers affects readers’ attachment to those documents, as virtual papers cannot hold significance as domestic objects in the same way. Meanwhile, authors have taken advantage of this fact and increased the resolution of their figures to bend conference page limits, and started publishing on arXiv to quickly attract attention to their new results. Again, we see the same themes: a pattern in computer vision was preceded by a similar shift in computer graphics, which has contributed to commodification. In this case, the conference publishing system has shifted from a means of scholarly communication to a means of evaluating research. Acceptance to a computer vision conference serves as a marker, not just of peer-reviewed technical correctness, but of sufficient novelty and significance to warrant high scores from reviewers, a “stamp of approval” for employment or promotion. In other words, improvements in writing tools have shifted the burden of evaluating new research work from the authors to the peer reviewers, placing significant labor burdens on the peer review process.
5. Discussion
Using a media archaeology approach, we have described the development of three aspects of the design of the contemporary computer vision research paper. First, we saw how teaser images, titles with acronyms and videos advertise the contribution of a paper, and how attention from arXiv and social media has become more important. Second, we saw how the results table was introduced for measuring the significance of a contribution, and become a ubiquitous part of the computer vision research paper. Finally, we looked at the transition from paper to PDF as a medium for disseminating research contributions and its new affordances for figure color and density. These trends have a key commonality: they make research papers easier to consume visually and more readily disposable.
These shifts showcase the changing material reality of academic publishing. The negligible reproduction cost of digital documents has reduced the labor required to communicate research. That has shifted the labor burden onto the attention of peer reviewers and online communities of scholars. The peer review process is now governed by self-imposed limiting factors, born of a desire by conferences to signal prestige through low acceptance rates. While we observe these changes in computer vision, they may soon appear in other disciplines as well.
When we view online attention as value-producing labor, we see a clear explanation for our visual trends. Authors who write more papers which are easy to understand at a glance, easy to promote on social media, more obviously novel and more significant in the eyes of reviewers attract more attention, generating increased benefit from the labor of their peers. This benefit comes in a variety of forms including citations, interest from future students and employees and increased progress on specific research problems. Authors seeking attention can develop design innovations, like more readable tables, but the attention economy ultimately privileges large industry labs. Discussion around the CVPR social media ban echoes these attempts to prevent competition between large industry labs and graduate students. But even social media bans cannot prevent the design of the paper itself from attracting attention. In reality, the value of that attention is great enough to motivate industry labs to invest in publishing. In addition to the marketing benefits, those companies can leverage scholarly attention to signal (Feldman and March, 1981) the innovation and pedigree needed to recruit the best computer vision researchers.
We believe this pattern exemplifies a broader trend of exploiting the attention of researchers. While authors derive prestige and career progress from online attention, the researchers who read those papers do not. Instead, paying attention to the state of the discipline is essential for doing career-progressing research of their own. While staying current has always been a part of academic work, the growth of the discipline puts pressure on it, increasingly treating researchers like a crowd of interchangeable peers whose attention is needed to make new research prestigious. This process resembles proletarianization333In Marxist theory, the process by which people are assimilated into the working class, causing alienation. (Bueno, 2017, Ch. 3), and may cause the “malaise” observed by Su and Crandall (Su and Crandall, 2021). Just as these trends in computer vision were foreshadowed by trends in computer graphics, similar trends may follow in other fields. In a recent paper examining paper titles and abstracts, Ken Hyland argues that an attention economy has begun to form widely throughout academic publishing (Hyland, 2023). We encourage future study of the attention economy to focus on both the labor performed by peers, as well as the sort of visual evidence we have discussed.
These problems become more egregious when considered alongside the much broader labor issues involving data labeling in computer vision (Paullada et al., 2021; Scheuerman et al., 2021; Denton et al., 2021). Workers have their intelligence commodified and exploited by researchers who are themselves responding to the commodification of their work. In this way, commodifying the labor of relatively privileged academics can exacerbate widespread exploitation of a less privileged workforce.
These problems also echo recent discussions regarding generative artificial intelligence. These technologies (some of which originate from CVPR, e.g. (Li et al., 2023)) show risk of producing online disinformation (Vykopal et al., 2023), displacing artists and writers (Jiang et al., 2023) and creating problems for future curators and historians regarding provenance and authenticity of historical images (Hazan, 2023). For example, as Jiang et al. discuss, the science fiction magazine Clarkesworld was forced to pause submissions due to excessive AI-generated submissions after ChatGPT went public in 2023 (Jiang et al., 2023). While authors submitting to computer vision conferences are not yet producing their papers using AI text generators, there are parallels regarding a deluge of content that places the burden of evaluation on the attention of reviewers. While bans on AI-generated content may slow the growth in the short-term, they do not resolve the fundamental commodification of attention, and may fail as new tools continue to blur the boundary between human and AI-generated writing.
The main implication is that ad hoc responses to symptoms of the attention economy, like banning discussion on social media or punishing low effort peer reviews will not change the pressures which are at play. Instead, computer vision is faced with a wicked problem (Buchanan, 1992), spanning peer review, hiring, promotion and tenure, in addition to academic discourse, which demands a larger-scale rethinking of the inherent tensions between facilitating scholarly communication and evaluating scholarship. Removing barriers to research communication appears to speed up the pace of research, but also changes disciplinary practices by forcing scholars into an arms race, competing for the attention of their peers.
Our inquiry also carries implications both for the design of the technologies which support publication and literature reviewing. As conferences continue to move beyond compatibility with paper, there is growing need for a LaTeX-compatible, digital-first document file format which gives authors control of the look and feel of their publication, but is machine-readable and supports accessibility tools. There is also a growing need for tools to help scholars collect and curate relevant literature. But rather than make this process as fast and frictionless as possible, we encourage work grounded in “slow design” (Odom et al., 2012) and “slow science” (Stengers, 2018) which attempts to recreate the preciousness, uniqueness and care that our participants ascribed to the papers they photocopied by hand. We also encourage, in line with the view of attention as labor, top conferences to consider compensating peer reviewers for the essential labor they perform.
Our work carries significant limitations. Our visual analysis is not systematic, and is likely biased towards visibly obvious trends, as well as trends which are present in published papers that have been digitized, ignoring the visual culture of posters, presentations, and rejected papers. As with any study of recent history, we cannot take a fully objective approach, as our personal experiences will skew our judgment. Our interview participants also only represent the views of senior researchers. A study of current students, junior faculty, and other younger authors and the “tricks” and “hacks” they use in their papers would be excellent future work.
Additionally, we have only scratched the surface of the media archaeology of the research paper, and there are numerous opportunities for future work. Alongside the patterns we analyze, sophisticated visual languages for system diagrams and renderings of visual features have developed. We encourage researchers to study these visual languages more systematically. We have also neglected to discuss the relationship between these sorts of readings of research papers with the much larger field of information science and bibliometrics (Ball, 2020). A more quantitative study which measures the relationships between visual features of research papers and the structures of citation graphs, or a modeling study like Weng et al. conduct on internet memes (Weng et al., 2012), may prove fruitful.
We worry that the stylistic conventions of the field may be constraining the types of ideas students and faculty are willing to pursue. While this is true for any set of disciplinary norms, the culture of benchmarks in computer vision fosters a particular mindset of technological determinism, where research becomes a matter of either finding the best performing model for an existing task or proposing a new task and constructing a benchmark dataset for it. Within this mindset, it seems inevitable that all possible visual perception tasks have computational solutions, and research is only a matter of finding them first.
We also worry that these stylistic conventions may be contributing to the safety and injustice issues which currently surround machine learning (Buolamwini and Gebru, 2018; Mehrabi et al., 2021). If authors are under pressure to publish more and faster, it is easiest to do that by overstating the significance of completed work. There are a range of behaviors which contribute to this problem: from neglecting to explore the limitations of a method, to only showing favorable evaluations, to outright results fabrication. While we do not claim that authors are engaging in such behaviors, we do observe an tendency to write papers like advertisements and only minimally discuss the downsides of advertised methods. These behaviors become problematic when engineers, stakeholders or researchers in other fields who are unfamiliar with the reality of conference publishing may take the claims in papers at face value, and put minimally tested systems into production based on a general trust of computation and data.
6. Conclusion
In this paper, we have examined the commodification of researcher attention through the visual evidence in computer vision research papers. These changes lead authors to include teaser images, and the presence of teasers leads researchers to develop methods which can more easily be presented as products, ready-made for readers to download, use, and cite. Second, the results table allows researchers to demonstrate the effectiveness of their neural network methods, but then the expectation of comparison to other methods forces other kinds of papers into the same evaluation paradigm. Finally, desire for color figures led conferences to adopt digital proceedings, which has led authors to create content which cannot be viewed on paper, necessitating screen-based reading. Through Bueno, we interpret attention as a necessary form of labor in academic publishing. These trends shift that labor onto the attention of researchers and especially peer reviewers. While better tools to support paper-reading (e.g. (Qian et al., 2019; Chang et al., 2023; Head et al., 2022)) can help streamline this process and reduce the burden on individual scholars, the problem will persist unless scholarly communities address the commodification of scholarly attention. Generally, we recommend that conference organizers approach this situation as a systemic problem, and investigate measures to avoid competition between authors, slow the submission and peer review process and search for more equitable and efficient ways of distributing attention and prestige.
Acknowledgements.
This work was supported in part by a NSF graduate research fellowship. The views expressed in this paper are those of the authors and do not necessarily reflect the views of the NSF or any other agencies.References
- (1)
- bou (2015) 2015. Bounty Paper Towels (Rubber Band Ball) Commercial. https://www.youtube.com/watch?v=-y30RsG6DJ8, screenshot taken by the authors..
- goo (2024) 2024. English Google Scholar metrics. https://scholar.google.com/citations?viewop=topvenues&hl=en
- pam (2024) 2024. List of Approved TC Motions. https://tc.computer.org/tcpami/tc-motions/
- Aydemir and Özsoy (2020) Cem Aydemir and SAMED Özsoy. 2020. Environmental impact of printing inks and printing process. Journal of Graphic Engineering and Design 2 (2020).
- Balan et al. (2007) Alexandru O. Balan, Leonid Sigal, Michael J. Black, James E. Davis, and Horst W. Haussecker. 2007. Detailed Human Shape and Pose from Images. In 2007 IEEE Conference on Computer Vision and Pattern Recognition. 1–8. https://doi.org/10.1109/CVPR.2007.383340
- Ball (2020) Rafael Ball (Ed.). 2020. Handbook Bibliometrics. De Gruyter Saur, Berlin, Boston. https://doi.org/doi:10.1515/9783110646610
- Bernard et al. (2016) Florian Bernard, Peter Gemmar, Frank Hertel, Jorge Goncalves, and Johan Thunberg. 2016. Linear Shape Deformation Models with Local Support Using Graph-Based Structured Matrix Factorisation. In 2016 IEEE Conference on Computer Vision and Pattern Recognition (CVPR). 5629–5638. https://doi.org/10.1109/CVPR.2016.607
- Berners-Lee (1989) Timothy J Berners-Lee. 1989. Information management: A proposal. Technical Report.
- Black (2021) Michael Black. 2021. PAMI-TC Meeting at CVPR 2021 Motion #4: Social-Media Limitation During Review. https://www.dropbox.com/s/mtrz5e5ezn4v94i/CVPR%202021%20Motion%204.pdf
- Blevis (2016) Eli Blevis. 2016. Being Photo-Visual in HCI and Design. In Proceedings of the 2016 ACM Conference on Designing Interactive Systems (DIS ’16). Association for Computing Machinery, Brisbane, QLD, Australia, 983–995. https://doi.org/10.1145/2901790.2901863
- Blevis et al. (2015) Eli Blevis, Sabrina Hauser, and William Odom. 2015. Sharing the Hidden Treasure in Pictorials. interactions 22, 3 (April 2015), 32–43. https://doi.org/10.1145/2755534
- Bogo et al. (2017) Federica Bogo, Javier Romero, Gerard Pons-Moll, and Michael J. Black. 2017. Dynamic FAUST: Registering Human Bodies in Motion. In 2017 IEEE Conference on Computer Vision and Pattern Recognition (CVPR). 5573–5582. https://doi.org/10.1109/CVPR.2017.591
- Brown and Cunningham (2007) Judy Brown and Steve Cunningham. 2007. A history of ACM Siggraph. Commun. ACM 50, 5 (2007), 54–61.
- Buchanan (1992) Richard Buchanan. 1992. Wicked problems in design thinking. Design issues 8, 2 (1992), 5–21.
- Bueno (2017) Claudio Celis Bueno. 2017. The Attention Economy: Labor time and power in cognitive capitalism. Rowman & Littlefield International.
- Buolamwini and Gebru (2018) Joy Buolamwini and Timnit Gebru. 2018. Gender shades: Intersectional accuracy disparities in commercial gender classification. In Conference on fairness, accountability and transparency. PMLR, 77–91.
- Buriak (2011) Jillian Buriak. 2011. Summarize your work in 100 milliseconds or less… the importance of the table of contents image. , 7687–7689 pages.
- Byun et al. (2021) Jaeseok Byun, Sungmin Cha, and Taesup Moon. 2021. Fbi-denoiser: Fast blind image denoiser for poisson-gaussian noise. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 5768–5777.
- Chan et al. (2021) Kelvin C.K. Chan, Xintao Wang, Xiangyu Xu, Jinwei Gu, and Chen Change Loy. 2021. GLEAN: Generative Latent Bank for Large-Factor Image Super-Resolution. In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 14240–14249. https://doi.org/10.1109/CVPR46437.2021.01402
- Chang et al. (2023) Joseph Chee Chang, Amy X Zhang, Jonathan Bragg, Andrew Head, Kyle Lo, Doug Downey, and Daniel S Weld. 2023. CiteSee: Augmenting Citations in Scientific Papers with Persistent and Personalized Historical Context. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–15.
- Chella et al. (1999) Antonio Chella, Vito Di Gesù, Ignazio Infantino, Daniela Intravaia, and Cesare Valenti. 1999. A cooperating strategy for objects recognition. In Shape, Contour and Grouping in Computer Vision. Springer, 264–274.
- Chen et al. (2021) Jian Chen, Meng Ling, Rui Li, Petra Isenberg, Tobias Isenberg, Michael Sedlmair, Torsten Möller, Robert S Laramee, Han-Wei Shen, Katharina Wünsche, et al. 2021. Vis30k: A collection of figures and tables from ieee visualization conference publications. IEEE Transactions on Visualization and Computer Graphics 27, 9 (2021), 3826–3833.
- Chen et al. (2017) Wen Chen, David J. Crandall, and Norman Makoto Su. 2017. Understanding the Aesthetic Evolution of Websites: Towards a Notion of Design Periods. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (CHI ’17). ACM, New York, NY, USA, 5976–5987. https://doi.org/10.1145/3025453.3025607
- Choi et al. (2021) Sungha Choi, Sanghun Jung, Huiwon Yun, Joanne T. Kim, Seungryong Kim, and Jaegul Choo. 2021. RobustNet: Improving Domain Generalization in Urban-Scene Segmentation via Instance Selective Whitening. In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 11575–11585. https://doi.org/10.1109/CVPR46437.2021.01141
- Dai et al. (2021) Zihang Dai, Hanxiao Liu, Quoc V Le, and Mingxing Tan. 2021. Coatnet: Marrying convolution and attention for all data sizes. Advances in Neural Information Processing Systems 34 (2021), 3965–3977.
- Davenport and Beck (2001) Thomas H Davenport and John C Beck. 2001. The attention economy. Ubiquity 2001, May (2001), 1–es.
- Davila et al. (2020) Kenny Davila, Srirangaraj Setlur, David Doermann, Bhargava Urala Kota, and Venu Govindaraju. 2020. Chart mining: A survey of methods for automated chart analysis. IEEE transactions on pattern analysis and machine intelligence 43, 11 (2020), 3799–3819.
- Deng et al. (2009) Jia Deng, Wei Dong, Richard Socher, Li-Jia Li, Kai Li, and Li Fei-Fei. 2009. Imagenet: A large-scale hierarchical image database. In 2009 IEEE conference on computer vision and pattern recognition. Ieee, 248–255.
- Denton et al. (2021) Emily Denton, Alex Hanna, Razvan Amironesei, Andrew Smart, and Hilary Nicole. 2021. On the genealogy of machine learning datasets: A critical history of ImageNet. Big Data & Society 8, 2 (2021), 20539517211035955.
- Divvala et al. (2014) Santosh K. Divvala, Ali Farhadi, and Carlos Guestrin. 2014. Learning Everything about Anything: Webly-Supervised Visual Concept Learning. In 2014 IEEE Conference on Computer Vision and Pattern Recognition. 3270–3277. https://doi.org/10.1109/CVPR.2014.412
- Doosti et al. (2020) Bardia Doosti, Shujon Naha, Majid Mirbagheri, and David J Crandall. 2020. Hope-net: A graph-based model for hand-object pose estimation. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 6608–6617.
- Dourish (2006) Paul Dourish. 2006. Implications for design. In Proceedings of the SIGCHI conference on Human Factors in computing systems. 541–550.
- Ethayarajh and Jurafsky (2020) Kawin Ethayarajh and Dan Jurafsky. 2020. Utility is in the eye of the user: A critique of NLP leaderboards. arXiv preprint arXiv:2009.13888 (2020).
- Fairclough (1993) Norman Fairclough. 1993. Critical discourse analysis and the marketization of public discourse: The universities. Discourse & society 4, 2 (1993), 133–168.
- Feldman and March (1981) Martha S Feldman and James G March. 1981. Information in Organizations as Signal and Symbol. Administrative Science Quarterly 26, 2 (1981), 171–186.
- Foucault (1969) Michel Foucault. 1969. The Archaeology of Knowledge. Vol. 1. London: Tavistock Publications. trans. Alan Sheridan Smith.
- Gaboury (2021) Jacob Gaboury. 2021. Image objects: An archaeology of computer graphics. MIT Press.
- Gao and Grauman (2021) Ruohan Gao and Kristen Grauman. 2021. Visualvoice: Audio-visual speech separation with cross-modal consistency. In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, 15490–15500.
- Garfield (1996) Eugene Garfield. 1996. What is the primordial reference for the phrase ‘publish or perish’. The Scientist 10, 12 (1996), 11.
- Geiger (2019) R Stuart Geiger. 2019. The Rise and Fall of the Note: Changing Paper Lengths in ACM CSCW, 2000-2018. Proceedings of the ACM on Human-Computer Interaction 3, CSCW (Nov. 2019), 222:1–222:10. https://doi.org/10.1145/3359324
- Gobert and Beaudouin-Lafon (2022) Camille Gobert and Michel Beaudouin-Lafon. 2022. i-LaTeX: Manipulating Transitional Representations between LaTeX Code and Generated Documents. In CHI Conference on Human Factors in Computing Systems. 1–16.
- Goldhaber (1997) Michael H Goldhaber. 1997. The attention economy and the net. First Monday (1997).
- Goree et al. (2022) Samuel Goree, David Crandall, and Norman Makoto Su. 2022. “It Was Really All About Books:” Speech-like Techno-Masculinity in the Rhetoric of Dot-Com Era Web Design Books. ACM Transactions on Computer-Human Interaction (2022).
- Goree et al. (2021) Samuel Goree, Bardia Doosti, David Crandall, and Norman Makoto Su. 2021. Investigating the homogenization of web design: A mixed-methods approach. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–14.
- Guo et al. (2021) Jianyuan Guo, Kai Han, Yunhe Wang, Han Wu, Xinghao Chen, Chunjing Xu, and Chang Xu. 2021. Distilling object detectors via decoupled features. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 2154–2164.
- Haber et al. (2014) Jonathan Haber, Miguel A Nacenta, and Sheelagh Carpendale. 2014. Paper vs. tablets: The effect of document media in co-located collaborative work. In Proceedings of the 2014 international working conference on advanced visual interfaces. 89–96.
- Hara et al. (2000) Shunsuke Hara, Nobuyuki Ohtake, Mika Higuchi, Noriko Miyazaki, Ayako Watanabe, Kanako Kusunoki, and Hiroshi Sato. 2000. MathBraille; a system to transform LATEX documents into Braille. ACM SIGCAPH Computers and the Physically Handicapped 66 (2000), 17–20.
- Hazan (2023) Susan Hazan. 2023. The Dance of the Doppelgängers: AI and the cultural heritage community. In Proceedings of EVA London 2023. BCS Learning & Development, 77–84.
- He et al. (2021) Yinan He, Bei Gan, Siyu Chen, Yichun Zhou, Guojun Yin, Luchuan Song, Lu Sheng, Jing Shao, and Ziwei Liu. 2021. Forgerynet: A versatile benchmark for comprehensive forgery analysis. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 4360–4369.
- Head et al. (2021) Andrew Head, Kyle Lo, Dongyeop Kang, Raymond Fok, Sam Skjonsberg, Daniel S Weld, and Marti A Hearst. 2021. Augmenting scientific papers with just-in-time, position-sensitive definitions of terms and symbols. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems. 1–18.
- Head et al. (2022) Andrew Head, Amber Xie, and Marti A Hearst. 2022. Math Augmentation: How Authors Enhance the Readability of Formulas using Novel Visual Design Practices. In CHI Conference on Human Factors in Computing Systems. 1–18.
- Hruby et al. (2022) Petr Hruby, Timothy Duff, Anton Leykin, and Tomas Pajdla. 2022. Learning to Solve Hard Minimal Problems. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 5532–5542.
- Huang (2018) Jia-Bin Huang. 2018. Deep paper gestalt. arXiv preprint arXiv:1812.08775 (2018).
- Huhtamo and Parikka (2011) Erkki Huhtamo and Jussi Parikka. 2011. Media archaeology: Approaches, applications, and implications. Univ of California Press. 3 pages.
- Huynh et al. (2018) Loc Huynh, Weikai Chen, Shunsuke Saito, Jun Xing, Koki Nagano, Andrew Jones, Paul Debevec, and Hao Li. 2018. Mesoscopic Facial Geometry Inference Using Deep Neural Networks. In 2018 IEEE/CVF Conference on Computer Vision and Pattern Recognition. 8407–8416. https://doi.org/10.1109/CVPR.2018.00877
- Hyland (2023) Ken Hyland. 2023. Academic publishing and the attention economy. Journal of English for Academic Purposes 64 (2023), 101253.
- Jiang et al. (2023) Harry H Jiang, Lauren Brown, Jessica Cheng, Mehtab Khan, Abhishek Gupta, Deja Workman, Alex Hanna, Johnathan Flowers, and Timnit Gebru. 2023. AI Art and its Impact on Artists. In Proceedings of the 2023 AAAI/ACM Conference on AI, Ethics, and Society. 363–374.
- Kasdorf (1998) Bill Kasdorf. 1998. SGML and PDF–Why We Need Both. Journal of Electronic Publishing 3, 4 (1998).
- Kocabas et al. (2020) Muhammed Kocabas, Nikos Athanasiou, and Michael J. Black. 2020. VIBE: Video Inference for Human Body Pose and Shape Estimation. In 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 5252–5262. https://doi.org/10.1109/CVPR42600.2020.00530
- Kong et al. (2022) Naejin Kong, Kiwoong Park, and Harshith Goka. 2022. Hole-robust Wireframe Detection. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 1636–1645.
- Kraut et al. (2002) Robert E Kraut, James Morris, Rahul Telang, Darrin Filer, Matt Cronin, and Shyam Sunder. 2002. Markets for attention: Will postage for email help?. In Proceedings of the 2002 ACM conference on Computer supported cooperative work. 206–215.
- Krizhevsky et al. (2012) Alex Krizhevsky, Ilya Sutskever, and Geoffrey E Hinton. 2012. ImageNet Classification with Deep Convolutional Neural Networks. In Advances in Neural Information Processing Systems, F. Pereira, C.J. Burges, L. Bottou, and K.Q. Weinberger (Eds.), Vol. 25. Curran Associates, Inc. https://proceedings.neurips.cc/paper/2012/file/c399862d3b9d6b76c8436e924a68c45b-Paper.pdf
- Kundu et al. (2016) Abhijit Kundu, Vibhav Vineet, and Vladlen Koltun. 2016. Feature Space Optimization for Semantic Video Segmentation. In 2016 IEEE Conference on Computer Vision and Pattern Recognition (CVPR). 3168–3175. https://doi.org/10.1109/CVPR.2016.345
- Latour (1987) Bruno Latour. 1987. Science in action: How to follow scientists and engineers through society. Harvard university press.
- Lee (2014) Icy Lee. 2014. Publish or perish: The myth and reality of academic publishing. Language teaching 47, 2 (2014), 250–261.
- Lee-Robbins and Adar (2022) Elsie Lee-Robbins and Eytan Adar. 2022. Affective Learning Objectives for Communicative Visualizations. (2022), 11.
- Li et al. (2021) Shichao Li, Zengqiang Yan, Hongyang Li, and Kwang-Ting Cheng. 2021. Exploring intermediate representation for monocular vehicle pose estimation. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 1873–1883.
- Li et al. (2023) Xiuyu Li, Yijiang Liu, Long Lian, Huanrui Yang, Zhen Dong, Daniel Kang, Shanghang Zhang, and Kurt Keutzer. 2023. Q-diffusion: Quantizing diffusion models. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 17535–17545.
- Lin et al. (2014) Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva Ramanan, Piotr Dollár, and C Lawrence Zitnick. 2014. Microsoft coco: Common objects in context. In European conference on computer vision. Springer, 740–755.
- Liu et al. (2021) Xiao-Chang Liu, Yong-Liang Yang, and Peter Hall. 2021. Learning to Warp for Style Transfer. In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 3701–3710. https://doi.org/10.1109/CVPR46437.2021.00370
- Luckiesh and Moss (1941) Matthew Luckiesh and Frank Kendall Moss. 1941. The effect of line-length on readability. Journal of Applied Psychology 25, 1 (1941), 67.
- Manovich (2002) Lev Manovich. 2002. The language of new media. MIT press.
- Manzoor et al. (2018) Ahtsham Manzoor, Murayyiam Parvez, Suleman Shahid, and Asim Karim. 2018. Assistive Debugging to Support Accessible Latex Based Document Authoring. In Proceedings of the 20th International ACM SIGACCESS Conference on Computers and Accessibility. 432–434.
- Mehrabi et al. (2021) Ninareh Mehrabi, Fred Morstatter, Nripsuta Saxena, Kristina Lerman, and Aram Galstyan. 2021. A survey on bias and fairness in machine learning. ACM computing surveys (CSUR) 54, 6 (2021), 1–35.
- Montfort et al. (2014) Nick Montfort, Patsy Baudoin, John Bell, Ian Bogost, and Jeremy Douglass. 2014. 10 PRINT CHR $(205.5+ RND (1));: GOTO 10. MIT Press.
- Odom et al. (2012) William Odom, Richard Banks, Abigail Durrant, David Kirk, and James Pierce. 2012. Slow technology: critical reflection and future directions. In Proceedings of the Designing Interactive Systems Conference. 816–817.
- Odom et al. (2009) William Odom, James Pierce, Erik Stolterman, and Eli Blevis. 2009. Understanding why we preserve some things and discard others in the context of interaction design. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 1053–1062.
- Ohme et al. (2011) Rafal Ohme, Michal Matukin, and Beata Pacula-Lesniak. 2011. Biometric measures for interactive advertising research. Journal of interactive advertising 11, 2 (2011), 60–72.
- Pasquinelli (2015) Matteo Pasquinelli. 2015. Italian operaismo and the information machine. Theory, Culture & Society 32, 3 (2015), 49–68.
- Patton (1990) Michael Quinn Patton. 1990. Qualitative Evaluation and Research Methods (second ed.). Thousand Oaks: Sage Publications, Inc.
- Paullada et al. (2021) Amandalynne Paullada, Inioluwa Deborah Raji, Emily M Bender, Emily Denton, and Alex Hanna. 2021. Data and its (dis) contents: A survey of dataset development and use in machine learning research. Patterns 2, 11 (2021), 100336.
- Pausch et al. (1996) Randy Pausch, Jon Snoddy, Robert Taylor, Scott Watson, and Eric Haseltine. 1996. Disney’s Aladdin: first steps toward storytelling in virtual reality. In Proceedings of the 23rd annual conference on Computer graphics and interactive techniques. 193–203.
- Pine and Liboiron (2015) Kathleen H Pine and Max Liboiron. 2015. The politics of measurement and action. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems. 3147–3156.
- Premachandran et al. (2014) Vittal Premachandran, Daniel Tarlow, and Dhruv Batra. 2014. Empirical Minimum Bayes Risk Prediction: How to Extract an Extra Few % Performance from Vision Models with Just Three More Parameters. In 2014 IEEE Conference on Computer Vision and Pattern Recognition. 1043–1050. https://doi.org/10.1109/CVPR.2014.137
- Qian et al. (2019) Xin Qian, Matt J Erhart, Aniket Kittur, Wayne G Lutters, and Joel Chan. 2019. Beyond iTunes for papers: Redefining the unit of interaction in literature review tools. In Conference Companion Publication of the 2019 on Computer Supported Cooperative Work and Social Computing. 341–346.
- Ramos and Concepcion (2020) Everly Ramos and Beatrice P Concepcion. 2020. Visual abstracts: redesigning the landscape of research dissemination. In Seminars in nephrology, Vol. 40. Elsevier, 291–297.
- Redmon et al. (2016) Joseph Redmon, Santosh Divvala, Ross Girshick, and Ali Farhadi. 2016. You only look once: Unified, real-time object detection. In Proceedings of the IEEE conference on computer vision and pattern recognition. 779–788.
- Redmon and Farhadi (2017) Joseph Redmon and Ali Farhadi. 2017. YOLO9000: better, faster, stronger. In Proceedings of the IEEE conference on computer vision and pattern recognition. 7263–7271.
- Rose (2016) Gillian Rose. 2016. Visual methodologies: An introduction to researching with visual materials. sage.
- Scheuerman et al. (2021) Morgan Klaus Scheuerman, Alex Hanna, and Emily Denton. 2021. Do datasets have politics? Disciplinary values in computer vision dataset development. Proceedings of the ACM on Human-Computer Interaction 5, CSCW2 (2021), 1–37.
- Sculley et al. (2018) David Sculley, Jasper Snoek, Alex Wiltschko, and Ali Rahimi. 2018. Winner’s curse? On pace, progress, and empirical rigor. (2018).
- Simon (1971) Herbert Simon. 1971. Designing organizations for an information-rich world. (1971), 37––52.
- Soto-Sanfiel et al. (2023) María T Soto-Sanfiel, Chin-Wen Chong, and José I Latorre. 2023. Hype in Science Communication: Exploring Scientists’ Attitudes and Practices in Quantum Physics. arXiv preprint arXiv:2311.07160 (2023).
- Stalling and Hege (1995) Detlev Stalling and Hans-Christian Hege. 1995. Fast and resolution independent line integral convolution. In Proceedings of the 22nd annual conference on Computer graphics and interactive techniques. 249–256.
- Stengers (2018) Isabelle Stengers. 2018. Another science is possible: A manifesto for slow science. John Wiley & Sons.
- Su and Crandall (2021) Norman Makoto Su and David J Crandall. 2021. The affective growth of computer vision. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 9291–9300.
- Tapaswi et al. (2016) Makarand Tapaswi, Yukun Zhu, Rainer Stiefelhagen, Antonio Torralba, Raquel Urtasun, and Sanja Fidler. 2016. MovieQA: Understanding Stories in Movies through Question-Answering. In 2016 IEEE Conference on Computer Vision and Pattern Recognition (CVPR). 4631–4640. https://doi.org/10.1109/CVPR.2016.501
- Taylor et al. (2015) Brian Taylor, Vasiliy Karasev, and Stefano Soattoc. 2015. Causal video object segmentation from persistence of occlusions. In 2015 IEEE Conference on Computer Vision and Pattern Recognition (CVPR). 4268–4276. https://doi.org/10.1109/CVPR.2015.7299055
- Triggs et al. (2000) Bill Triggs, Andrew Zisserman, and Richard Szeliski. 2000. Vision Algorithms: Theory and Practice: International Workshop on Vision Algorithms Corfu, Greece, September 21–22, 1999 Proceedings. Springer.
- Ummenhofer et al. (2017) Benjamin Ummenhofer, Huizhong Zhou, Jonas Uhrig, Nikolaus Mayer, Eddy Ilg, Alexey Dosovitskiy, and Thomas Brox. 2017. Demon: Depth and motion network for learning monocular stereo. In Proceedings of the IEEE conference on computer vision and pattern recognition. 5038–5047.
- van Koningsbruggen and Hornecker (2021) Rosa van Koningsbruggen and Eva Hornecker. 2021. “It’s Just a Graph” – The Effect of Post-Hoc Rationalisation on InfoVis Evaluation. In Creativity and Cognition (C&C ’21). Association for Computing Machinery, New York, NY, USA, 1–10. https://doi.org/10.1145/3450741.3465257
- Vardi (2009) Moshe Y. Vardi. 2009. Conferences vs. Journals in Computing Research. Commun. ACM 52, 5 (May 2009), 5–5. https://doi.org/10.1145/1506409.1506410
- Von Bearnensquash (2010) Carven Von Bearnensquash. 2010. Paper gestalt. Secret Proceedings of Computer Vision and Pattern Recognition (CVPR) (2010).
- Vykopal et al. (2023) Ivan Vykopal, Matúš Pikuliak, Ivan Srba, Robert Moro, Dominik Macko, and Maria Bielikova. 2023. Disinformation Capabilities of Large Language Models. arXiv preprint arXiv:2311.08838 (2023).
- Wallace et al. (2017) James R. Wallace, Saba Oji, and Craig Anslow. 2017. Technologies, Methods, and Values: Changes in Empirical Research at CSCW 1990 - 2015. Proceedings of the ACM on Human-Computer Interaction 1, CSCW (Dec. 2017), 106:1–106:18. https://doi.org/10.1145/3134741
- Wang et al. (2021) Shaofei Wang, Andreas Geiger, and Siyu Tang. 2021. Locally Aware Piecewise Transformation Fields for 3D Human Mesh Registration. In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 7635–7644. https://doi.org/10.1109/CVPR46437.2021.00755
- Weinzaepfel et al. (2011) Philippe Weinzaepfel, Hervé Jégou, and Patrick Pérez. 2011. Reconstructing an image from its local descriptors. In CVPR 2011. 337–344. https://doi.org/10.1109/CVPR.2011.5995616
- Weng et al. (2012) Lilian Weng, Alessandro Flammini, Alessandro Vespignani, and Fillipo Menczer. 2012. Competition among memes in a world with limited attention. Scientific reports 2, 1 (2012), 335.
- Wilkinson and Light (2011) Krista M Wilkinson and Janice Light. 2011. Preliminary investigation of visual attention to human figures in photographs: Potential considerations for the design of aided AAC visual scene displays. (2011).
- Willett et al. (2021) Wesley Willett, Bon Adriel Aseniero, Sheelagh Carpendale, Pierre Dragicevic, Yvonne Jansen, Lora Oehlberg, and Petra Isenberg. 2021. Superpowers as inspiration for visualization. IEEE TVCG 2021 (2021).
- Wu et al. (2021) Aoyu Wu, Yun Wang, Xinhuan Shu, Dominik Moritz, Weiwei Cui, Haidong Zhang, Dongmei Zhang, and Huamin Qu. 2021. Ai4vis: Survey on artificial intelligence approaches for data visualization. IEEE Transactions on Visualization and Computer Graphics (2021).
- Yardi et al. (2009) Sarita Yardi, Scott A Golder, and Michael J Brzozowski. 2009. Blogging at work and the corporate attention economy. In Proceedings of the SIGCHI conference on human factors in computing systems. 2071–2080.
- Zhang et al. (2017) Chao Zhang, Sergi Pujades, Michael Black, and Gerard Pons-Moll. 2017. Detailed, Accurate, Human Shape Estimation from Clothed 3D Scan Sequences. In 2017 IEEE Conference on Computer Vision and Pattern Recognition (CVPR). 5484–5493. https://doi.org/10.1109/CVPR.2017.582
- Zheng et al. (2021) Mianlun Zheng, Yi Zhou, Duygu Ceylan, and Jernej Barbič. 2021. A Deep Emulator for Secondary Motion of 3D Characters. In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). 5928–5936. https://doi.org/10.1109/CVPR46437.2021.00587

### v3

@@
 Specifically, we study these changes in computer vision over the past decade, as the deep learning revolution has driven unprecedented growth in the discipline.
-We ground our investigation through interviews with veteran researchers spanning computer vision, graphics and visualization.
-Our analysis focuses on the research attention economy: how research paper elements contribute towards advertising, measuring and disseminating an increasingly commodified “contribution.” Through this work, we seek to motivate future discussion surrounding the design of both the research paper itself as well as the larger sociotechnical research publishing system, including tools for finding, reading and writing research papers.
+We ground our investigation through interviews with veteran researchers spanning computer vision, graphics, and visualization.
+Our analysis focuses on the research attention economy: how research paper elements contribute towards advertising, measuring, and disseminating an increasingly commodified “contribution.” Through this work, we seek to motivate future discussion surrounding the design of both the research paper itself as well as the larger sociotechnical research publishing system, including tools for finding, reading, and writing research papers.
 Several teaser images from computer vision research papers, including elements like various 3D models of human faces and bodies, images with bounding boxes and distorted images.
@@
 In this paper, we study the changes to computing papers in one area—computer vision—over the past decade.
-While changes occur over time in the research papers of every discipline, computer vision is a particularly information-rich site (Patton, 1990, p.242) for understanding the paper as a media artifact both because of its inherently visual nature and its unprecedented recent growth.
+While changes occur over time in the research papers of every discipline, computer vision is a particularly information-rich site (Patton, 1990, p.
+242) for understanding the paper as a media artifact both because of its inherently visual nature and its unprecedented recent growth.
 Over the past decade, the “deep learning revolution” has transformed several fields of computer science altering both the fields themselves and the way researchers and practitioners feel about them (Su and Crandall, 2021).
@@
 To understand its commodification, we turn to the post-Marxist analysis of Claudio Celis Bueno (Bueno, 2017).
-Bueno frames online attention, not just as a scarce resource, but as a form of labor—human activity which has the potential to generate surplus value.
+Bueno frames online attention not just as a scarce resource, but as a form of labor—human activity which has the potential to generate surplus value.
 We claim that this theory of attention as labor applies to the attention paid to academic research as well, and helps to explain recent changes to computer vision.
@@
 Perhaps the biggest hint that the design of computer vision papers has changed with the rapid changes in its disciplinary culture is scholarly humor.
-In the 2010 computer vision satire paper “Paper Gestalt” (Von Bearnensquash, 2010), the pseudonymous authors “take the simple intuition that the quality of a paper can be estimated by merely glancing through the general layout, and use this intuition to build a system that employs basic computer vision techniques to predict if the paper should be accepted or rejected,” (Von Bearnensquash, 2010) and suggest that this system might replace the peer review process.
+In the 2010 computer vision satire paper “Paper Gestalt” (Von Bearnensquash, 2010), the pseudonymous authors “take the simple intuition that the quality of a paper can be estimated by merely glancing through the general layout, and use this intuition to build a system that employs basic computer vision techniques to predict if the paper should be accepted or rejected” (Von Bearnensquash, 2010), and suggest that this system might replace the peer review process.
 The visual design of the paper parodies the qualities it observes, including unnecessary complex equations and long algorithms.
@@
 To better understand these developments in context, we also report results of interviews with twelve veteran researchers whose work spans computer vision, graphics, and visualization.
-Our approach focuses on elements of the research paper and larger publication system based on their roles in advertising, measuring, and disseminating an increasingly commodified “contribution.” We find that as academic discourse has moved online, the limiting factor on publishing has shifted from printing costs to the attention of other researchers, which has changed both the culture of computer vision, as well as the design of its research papers.
-While this finding echoes existing discourse surrounding “publish or perish” (Garfield, 1996; Lee, 2014), the “marketization” of academic discourse (Fairclough, 1993), use of “hype” in science communication (Soto-Sanfiel et al., 2023) or promotional qualities in titles and abstracts (Hyland, 2023), we emphasize the novelty of our approach.
+Our approach focuses on elements of the research paper and larger publication system based on their roles in advertising, measuring, and disseminating an increasingly commodified “contribution.” We find that as academic discourse has moved online, the limiting factor on publishing has shifted from printing costs to the attention of other researchers, which has changed both the culture of computer vision as well as the design of its research papers.
+While this finding echoes existing discourse surrounding “publish or perish” (Garfield, 1996; Lee, 2014), the “marketization” of academic discourse (Fairclough, 1993), use of “hype” in science communication (Soto-Sanfiel et al., 2023), or promotional qualities in titles and abstracts (Hyland, 2023), we emphasize the novelty of our approach.
 To the best of our knowledge ours is the first connection between these discussions of scholarly discourse and the designed visual artifact of the research paper.
@@
 Manzoor et al.
-develop a LaTeX editor extension to improve the accessibility of LaTeX for authors with visual impairments (Manzoor et al., 2018) and Hara et al.
+develop a LaTeX editor extension to improve the accessibility of LaTeX for authors with visual impairments (Manzoor et al., 2018), and Hara et al.
 develop a system for generating Braille documents from mathematical expressions written in LaTeX (Hara et al., 2000).
@@
 Motivated by the dangers of algorithmic discrimination and safety concerns in systems relying on computer vision algorithms, several recent papers have studied the culture of the computer vision research community and its understanding of data and truth, often using research papers as texts for analysis.
-Su and Crandall study the emotional state of the computer vision community, finding that the deep learning revolution and subsequent growth has had a profound effect on its culture, and leading both to excitement regarding progress as well as to isolation and malaise (Su and Crandall, 2021).
+Su and Crandall study the emotional state of the computer vision community, finding that the deep learning revolution and subsequent growth have had a profound effect on its culture, leading both to excitement as well as isolation and malaise (Su and Crandall, 2021).
 Denton et al.
@@
 and Ethayarajh and Jurafsky critique the concept of benchmarks and leaderboards in machine learning, arguing in different ways that steadily increasing scores do not always correspond to progress (Sculley et al., 2018; Ethayarajh and Jurafsky, 2020).
-For a more comprehensive survey please see Paullada et al.
+For a more comprehensive survey, please see Paullada et al.
 (Paullada et al., 2021).
@@
 Our inquiry began through visual analysis of historical computer vision research papers.
-We found the changes to the visual design of these papers surprising, particularly the increasing prevalence of highly complex figures, and decided to investigate further by conducting both semi-structured interviews with researchers who had been active in computer vision, graphics, and visualization for several decades, and computational analysis of a corpus of research papers published between 2013 and 2021 in the IEEE Conference on Computer Vision and Pattern Recognition—the highest ranked publication venue in computer science and third highest in all of science, according to Google Scholar (goo, 2024).
+We were surprised by visual design changes across these papers, particularly the increasing prevalence of highly complex figures, and decided to investigate further by conducting semi-structured interviews with researchers who had been active in computer vision, graphics, and visualization for several decades, as well as computational analysis of a corpus of research papers published between 2013 and 2021 in the IEEE/CVF Conference on Computer Vision and Pattern Recognition—the highest-ranked publication venue in computer science and second highest in all of science, according to Google Scholar as of August 2024 (goo, 2024).
 We study this conference, rather than similar conferences in machine learning or natural language processing, because of the importance of visual presentation in computer vision.
-Along the way, our visual analysis, interviews and computational findings condensed into a media archaeology approach.
+Along the way, our visual analysis, interviews, and computational findings coalesced into a media archaeology approach.
 3.1. Media Archaeology
 Inspired by recent work on the history of computer graphics (Gaboury, 2021), we examine historical research papers through the lens of media archaeology.
-By “archaeology,” this method refers to Foucault’s concept of archaeology as the “uncovering of the archive” (Foucault, 1969, p.131), rather than physical excavation.
+“Archaeology” here refers to Foucault’s concept of archaeology as the “uncovering of the archive” (Foucault, 1969, p.
+131), rather than physical excavation.
 Media archaeology, as defined by Huhtamo and Parikka, “rummages textual, visual and auditory archives as well as collections of artifacts, emphasizing both the discursive and material manifestations of culture” (Huhtamo and Parikka, 2011, p.
 3).
-In other words, we study digital media objects as components of a media culture; a system of practices and meanings which structure our interpretation (Rose, 2016).
+In other words, we study digital media objects as components of a media culture: a system of practices and meanings which structure our interpretation (Rose, 2016).
 We also consider their digital materiality as physical patterns of bits on hard drives, produced through the labor of researchers.
-By drawing attention to material culture, media archaeology is in dialogue with theories of historical materialism, which consider economic factors, production and consumption of goods and services, as the driving force of history, rather than revolutionary ideas or progress narratives (Gaboury, 2021, p.
+By drawing attention to material culture, media archaeology is in dialogue with theories of historical materialism, which consider economic factors—production and consumption of goods and services—as the driving force of history, rather than revolutionary ideas or progress narratives (Gaboury, 2021, p.
 5).
@@
 The goal of media archaeology is to dissect these inferences and identify how their layers of meaning developed over time.
-In addition to our computational analysis and detailed study of the CVPR proceedings, we “rummaged” research papers from the IEEE, Computer Vision Foundation, and ACM digital archives, including the IEEE International Conference on Computer Vision (ICCV), European Conference on Computer Vision (ECCV), IEEE Transactions on Pattern Analysis and Machine Intelligence (PAMI), and the ACM Speical Interest Group on Graphics and Interactive Techniques conference (SIGGRAPH).
+In addition to our computational analysis and detailed study of the CVPR proceedings, we “rummaged” research papers from the IEEE, Computer Vision Foundation, and ACM digital archives, including the IEEE International Conference on Computer Vision (ICCV), European Conference on Computer Vision (ECCV), IEEE Transactions on Pattern Analysis and Machine Intelligence (PAMI), and the ACM Special Interest Group on Graphics and Interactive Techniques conference (SIGGRAPH).
 We examined ICCV, ECCV, and PAMI in addition to CVPR because they are the four most influential computer vision venues according to Google Scholar,111https://web.archive.org/web/20220709182958/https://scholar.google.com/citations?view_op=top_venues&hl=en&vq=eng_computervisionpatternrecognition and SIGGRAPH due to its historical significance and prevalence in our interviews.
@@
 Participants are listed by years of industry and academic experience in Table 1.
-This sample is highly non-representative, showing both survival bias based on who remains in the field over a long period of time and selection bias based on who agreed to an interview, so it does not serve as the source of our claims.
+This sample is highly non-representative, showing both survival bias based on who remains in the field over a long period of time and selection bias based on who agreed to an interview, so it does not serve as the sole source of our claims.
 Instead, it provides important social context to our visual analysis.
@@
 After transcribing the interviews, we engaged in a process of iterative memoing, integrated into our weekly meetings where we compared examples from historical research papers and interview excerpts which shared common themes.
-We focused our analysis on identifying patterns from the context that participants provided while telling stories about their writing, as well how participants thought about their papers in relation to changing disciplinary practices, rather than simply analyzing the papers participants discussed, or discussing our participants’ analysis of their own work.
+We focused our analysis on identifying patterns from the context that participants provided while telling stories about their writing, as well how participants thought about their papers in relation to changing disciplinary practices, rather than simply analyzing the papers participants discussed or discussing our participants’ analysis of their own work.
 | Participant | Discipline | Years Industry | Years Academia |
@@
 Specifically, we were interested in whether teaser images, figures and tables were becoming more prevalent in CVPR papers over time, and whether more titles were following a particular format containing an acronym followed by a colon.
-To answer these questions, we collected open access PDF files from thecvf.com published in CVPR 2013 to 2021 and article metadata from ieeexplore.ieee.org for CVPR 1992 to 2020.
+To answer these questions, we collected PDFs of papers from the Computer Vision Foundation’s Open Access Repository (thecvf.com) published in CVPR 2013 to 2021 and article metadata from IEEE Xplore (ieeexplore.ieee.org) for CVPR 1992 to 2020.
 We manually inspected PDFs from thecvf.com to count teaser images, then used the Linux pdftotext tool combined with regular expressions to count figures and tables.
-From the IEEExplore metadata, we used regular expressions to parse paper titles, treating a word with more than two capital letters as an acronym, and an acronym as unique if it only appears once in a given year of data.
+From the Xplore metadata, we used regular expressions to parse paper titles, treating a word with more than two capital letters as an acronym, and an acronym as unique if it only appears once in a given year of data.
 4. Findings
@@
 3 passes, 3.5 sec elapsed time, 1 stp caramel, 0.5 psi.”
-Through our analysis, we found that several elements of the contemporary computer vision research paper serve as material traces of the disciplinary change which took place over the course of the 2010s.
+Through our analysis, we found that several elements of the contemporary computer vision research paper serve as material traces of the disciplinary change which took place during the 2010s.
 We find that papers are written to promote a specific contribution which is increasingly commodified.
-By “contribution,” we mean the model, algorithm, method, dataset or other system that the paper offers to readers.
+By “contribution,” we mean the model, algorithm, method, dataset, or other system that the paper offers to readers.
 By commodified, we mean reduced to its exchange value in terms of the attention it generates.
@@
 This process can take the form of feedback, where authors improve their work based on an early online response, or simply where early viewers find a paper interesting and share it, judging it as valuable, putting it in front of more eyes and increasing the chance that it will reach the right audience.
-Eventually, that attention translates into citations, jobs, students, grants and other benefits of academic prestige.
-We study three aspects of contemporary research papers: the teaser image, the results table and the high resolution figure.
+Eventually, that attention translates into citations, jobs, students, grants, and other benefits of academic prestige.
+We study three aspects of contemporary research papers: the teaser image, the results table, and the high resolution figure.
 We have grouped our results into thematic sections surrounding these three elements in relation to the concept of a paper’s contribution.
@@
 4.1.1. Teaser images
-The first element of research paper design that we examine is the “teaser image,” a large first-page figure which summarizes the paper.
+The “teaser image” is a large first-page figure which summarizes the paper.
 Several examples of teaser images are shown in Figures 1 and 2.
-These figures are functionally similar to the trend of “visual abstracts” (Ramos and Concepcion, 2020) and “table of contents images” (Buriak, 2011) in the biomedical sciences, however unlike these forms which are graphical summaries of the text, teaser images are part of the main paper and are usually a visualized system output that the authors want to show.
-Teaser images have been steadily gaining popularity within CVPR, as shown in Figure 3 left.
-For this analysis, we defined the teaser image as a first-page figure which covers the entire width of the page, not just a column.
+These figures are functionally similar to the trend of “visual abstracts” (Ramos and Concepcion, 2020) and “table of contents images” (Buriak, 2011) in the biomedical sciences; however unlike these forms which are graphical summaries of the text, teaser images are part of the main paper and are usually a visualized system output that the authors want to show.
+Teaser images have been steadily gaining popularity at CVPR, as shown in Figure 3 (left).
+For this analysis, we define the teaser image as a first-page figure which covers the entire width of the page, not just a column.
 Two subfigures.
@@
 P12 ascribes the teaser image to the famous graphics researcher Randy Pausch: “Randy Pausch…takes credit for putting the teaser image in SIGGRAPH papers…he claimed that he did a paper where half the first page was a teaser image and after that that became the norm where people started always putting these images.” The paper in question is pictured in Figure 2 (a).
-While Pausch did not invent the large first-page figure (Tim Berners-Lee famously used one in his 1989 proposal for the world wide web, for example (Berners-Lee, 1989)), he did publish the first paper in SIGGRAPH with this layout, and it quickly became popular.
-P2 and P10 point to the quick adoption of teaser images, which even became institutionalized in the SIGGRAPH template, which is now the template for all ACM conferences (Figure 2 (c).
+While Pausch did not invent the large first-page figure (Tim Berners-Lee famously used one in his 1989 proposal for the world wide web (Berners-Lee, 1989)), he did publish the first paper in SIGGRAPH with this layout, and it quickly became popular.
+P2 and P10 point to the quick adoption of teaser images, which even became institutionalized in the SIGGRAPH template and later the template for all ACM conferences (Figure 2 (c)).
 P12 defines the teaser as a visualization of either a result or a system and explains that it has spread to many other conferences because of the way it attracts attention: “It’s a trailer.
@@
 P2 echoes that sentiment: “They just made the papers look good!
-I mean, It’s much more memorable and there are some papers still today that I don’t remember the title, but you see the picture and you’re like, oh yeah, that’s the Randy Pausch paper on the VR for whatever, right?
-So yeah there’s a few of those that are just like, really iconic first pagers.” P2 is referencing the same Pausch paper as P12 (Pausch et al., 1996), highlighting its memorability.
+I mean, it’s much more memorable and there are some papers still today that I don’t remember the title, but you see the picture and you’re like, ‘Oh yeah, that’s the Randy Pausch paper on the VR for whatever, right?’ So yeah there’s a few of those that are just like, really iconic first pagers.” P2 is referencing the same Pausch paper as P12 (Pausch et al., 1996), highlighting its memorability.
 The teaser image is a trailer, a hook; it advertises the paper to the potential reader.
-The authors want to promote their paper and showcase the best results they can because the sheer number of and easy access to papers has made it harder to stand out.
+The authors want to promote their paper and showcase the best results they can because the sheer number and easy access to papers has made it harder to stand out.
 The visual organization of these figures echoes that theme.
@@
 We observe a similar advertising quality in titles.
-Figure 3 right shows the rise in popularity of a particular title construction where an acronym, which is usually the name of a model, is followed by a colon: “HOPE-Net: A Graph-Based Model for Hand-Object Pose Estimation” (Doosti et al., 2020) or “DeMoN: Depth and Motion Network for Learning Monocular Stereo” (Ummenhofer et al., 2017).
+Figure 3 (right) shows the rise in popularity of a particular title construction where an acronym, which is usually the name of a model, is followed by a colon: “HOPE-Net: A Graph-Based Model for Hand-Object Pose Estimation” (Doosti et al., 2020) or “DeMoN: Depth and Motion Network for Learning Monocular Stereo” (Ummenhofer et al., 2017).
 These names both signal that the paper proposes a new model and brands the paper with a short, memorable name which is often a cultural reference or clever pun.
@@
 Why would many authors begin to engage in these teaser figure and title practices over the course of the 2010s?
-We find that it has resulted from a change in medium through which research reaches its audience, from conference and journal proceedings onto social media and preprint websites.
-P4 describes how his students have strategies to get their papers noticed on the preprint site arXiv: “if they take your submission on Thursday and it goes up on the arXiv most recent publications list it’s up for all of Friday, Saturday, and Sunday…so it gets more days of exposure…It improves your odds that somebody will notice your thing.” P8 describes the pressure on her students to promote their work: “social media, like promoting one’s research has become such a big thing and I think students…are realizing, oh it’s not enough for me to, y’know, come up with a paper, post it on arXiv, get it accepted.
-I also need to tweet about it.
-It’s frankly quite exhausting…[before] the only way you promote your paper is it shows up at the conference and hopefully, y’know, some famous computer vision researchers will come up and look.” She talks about how her advisor would bring his friends over to her poster, and that was all the advertising she needed to get noticed, but that strategy no longer works as well in the modern crowded field of computer vision.
+One explanation is a change in medium through which research reaches its audience, from conference and journal proceedings onto social media and preprint websites.
+P4 describes how his students have strategies to get their papers noticed on the preprint site arXiv: “if they take your submission on Thursday and it goes up on the arXiv most recent publications list, it’s up for all of Friday, Saturday, and Sunday…so it gets more days of exposure…It improves your odds that somebody will notice your thing.” P8 describes the pressure on her students to promote their work: “social media, like promoting one’s research has become such a big thing and I think students…are realizing, ‘Oh it’s not enough for me to, y’know, come up with a paper, post it on arXiv, get it accepted.
+I also need to tweet about it.’ It’s frankly quite exhausting…[before] the only way you promote your paper is it shows up at the conference and hopefully, y’know, some famous computer vision researchers will come up and look.” She talks about how her advisor would bring his friends over to her poster, and that was all the advertising she needed to get noticed, but that strategy no longer works as well in the modern crowded field of computer vision.
 Again this pattern resembles one from computer graphics.
@@
 He switched fields from computational complexity to computer vision because there were more active researchers, and he could not tell if his work was having an impact in a small field.
-P9 tells the story of how attention from a senior scholar improved his relationship with his advisor: “that email from [scholar], that’s the beginning of our of my great relationship with my advisor was that one little email, because within five minutes [advisor] was up in the office, ‘Oh, you got an email from [scholar]?
+P9 tells the story of how attention from a senior scholar improved his relationship with his advisor: “that email from [scholar], that’s the beginning of my great relationship with my advisor was that one little email, because within five minutes [advisor] was up in the office, ‘Oh, you got an email from [scholar]?
 Very good.
 Very good.
-How about a coffee?’ ” Notice that the career benefits resulted, not from the attention itself, but from the record of that attention in the form of an email from a senior scholar.
+How about a coffee?’ ” Notice that the career benefits resulted not from the attention itself, but from the record of that attention in the form of an email from a senior scholar.
 While the relationship between attention and prestige is not new, the unprecedented growth of computer vision over the past decade places pressure on the attention of senior researchers.
@@
 Some of the proposed ideas are recorded in the passed motions of the annual meeting of the IEEE pattern analysis and machine intelligence technical committee (PAMI-TC), the governing body of the computer vision research community.
-In 2021-2022, motions changed the rules regarding double submission to reduce burden on reviewers, introduced punitive measures for missing or poor quality peer reviews from authors and banned discussion of submitted papers on social media to avoid influencing peer reviewers, the last of which was repealed in 2023 (pam, 2024; Black, 2021).
+In 2021–2022, motions changed the rules regarding double submission to reduce burden on reviewers, introduced punitive measures for missing or poor quality peer reviews from authors, and banned discussion of submitted papers on social media to avoid influencing peer reviewers (the last of which was repealed in 2023) (pam, 2024; Black, 2021).
 The rationale statement for the social media ban specifically highlights attention: “Groups with large followings and the resources to mount visible social media promotions received significant attention for work that is under review.
-Reviewers are exposed to this work and the attention it receives can bias their judgment—if so many people on social media are excited, mustn’t it be good?” (Black, 2021) In addition to highlighting the valorizing role of attention, this statement showcases the view of peer review as an unbiased process for evaluating the quality of research.
+Reviewers are exposed to this work and the attention it receives can bias their judgment—if so many people on social media are excited, mustn’t it be good?” (Black, 2021).
+In addition to highlighting the valorizing role of attention, this statement showcases the view of peer review as an unbiased process for evaluating the quality of research.
 We see two additional key themes starting to emerge.
@@
 But all of this was not the case a few decades ago.
-The vast majority of computer vision papers from the 1980s and 1990s rely on mathematical arguments based on pinhole camera geometry and do not contain any quantitative results.
+The vast majority of computer vision papers from the 1980s and 1990s rely on mathematical arguments based on, for example, pinhole camera geometry and do not contain any quantitative results.
 Empirical evaluation, if included at all, was primarily qualitative, in the form of figures showing sample results.
@@
 (f) shows a table with many rows split into sections with double horizontal lines, two numbers in bold and several cells with infinity symbols.
-The publication of Krizhevsky, Sutskever and Hinton’s “ImageNet Classification with Deep Convolutional Neural Networks” in 2012 (Krizhevsky et al., 2012) marks a turning point for empirical evaluation.
+The publication of Krizhevsky, Sutskever, and Hinton’s “ImageNet Classification with Deep Convolutional Neural Networks” in 2012 (Krizhevsky et al., 2012) marks a turning point for empirical evaluation.
 This paper is historically significant for setting off the deep learning revolution, and its design and writing served as a foundation for the thousands of deep learning-based computer vision papers that followed.
@@
 Neural network papers are obligated to use empirical evaluation, as there are insufficient theoretical guarantees for these models and they are difficult to evaluate otherwise.
-Over the following years, many papers followed Krizhevsky’s lead, showing that deep convolutional neural networks outperform existing methods on other central problems like object detection and semantic segmentation.
+Over the following years, many papers followed Krizhevsky et al.’s lead, showing that deep convolutional neural networks outperform existing methods on other central problems like object detection and semantic segmentation.
 We can see a corresponding increase in both the average number of tables per paper, as well as the fraction of papers containing at least one table in Figure 4.
 While the prevalence of figures has remained relatively constant, tables have become significantly more common.
-While only 75% of CVPR 2013 papers had a table, 95% of CVPR 2021 papers did, and the average number of tables per paper has doubled from 2 to 4.
-While not all of these tables are the kind of results table we’re discussing, the change is striking.
+While only 75% of CVPR 2013 papers had a table, 95% of CVPR 2021 papers did, and the average number of tables per paper has doubled from two to four.
+While not all of these tables are the kind of results table we are discussing, the change is striking.
 Like teaser images, however, this style of table arises in computer graphics before entering computer vision; see Figure 5 (a) for an early example.
@@
 Even highly geometric papers, like the example in Figure 5 (f), now involve empirical evaluation.
-More broadly, in computer vision there is widespread assumption that research inherently involves competition between technical methods.
+More broadly, in computer vision there is a widespread assumption that research inherently involves competition between technical methods.
 Several of our participants described academic publication using free market metaphors.
@@
 P4 described vision as “very industrial” and gave examples of techniques that students use to optimize their arXiv submissions to reach as many eyes as possible.
-While these tables are buried deep within the paper, it may seem like their effect on attention is minimal.
+With these tables buried deep within the paper, it may seem like their effect on attention is minimal.
 However, a competitive benchmark result attracts attention on its own.
 Once a researcher has the lead, other papers must compare against their benchmark and cite them.
-P6 compares benchmarks to arcade game leaderboards: “If it’s an established benchmark, there’s like there’s somebody who has the lead right?
+P6 compares benchmarks to arcade game leaderboards: “If it’s an established benchmark… there’s somebody who has the lead right?
 Like you would have on like a classic video game arcade machine, right?
@@
 That’s exactly what they’re doing, and that’s frustrating.” P5 and P8 advise their students to avoid working on research which forces them into competition with large companies.
-P5: “I tell them, don’t work on problems that, you know, a lot of people are working on right now, you can’t possibly compete with Facebook, Google, Amazon because you’re not as computing heavy.” While students could still develop innovative ideas for those problems, they likely would not beat the benchmarks that large companies have set without leveraging similar computing power and would not attract attention.
+P5 explains, “I tell them, don’t work on problems that, you know, a lot of people are working on right now, you can’t possibly compete with Facebook, Google, Amazon because you’re not as computing heavy.” While students could still develop innovative ideas for those problems, they likely would not beat the benchmarks that large companies have set without leveraging similar computing power and would not attract attention.
+Because students will not be able to compete with large companies for attention on these problems, they are not worth studying.
 In summary, the development of the results table shows the transition of computer vision from an applied mathematical discipline to a quantitative-empirical one.
 As vision started relying on empirical evaluation, it adopted a style of results table which was used for runtime benchmarking in computer graphics, and as evaluation benchmarks became established for vision problems, these tables grew in size and importance.
-Today, competition on benchmarks is the organizing principle of the discipline, new methods must demonstrate that they are empirically more effective than existing methods to be accepted, and the results table is an essential part of the paper.
+Today, competition on benchmarks is the organizing principle of the discipline; new methods must demonstrate that they are empirically more effective than existing methods to be accepted, and the results table is an essential part of the paper.
 This element showcases several of the same themes as teaser images: influence from computer graphics, and commodification, in the form of measurable improvements over prior work.
@@
 The final element we explore is the Portable Document Format (PDF) itself, and the way that authors engage with it as a medium.
-In this section, to avoid confusion, we use “research paper” to refer to printed papers and “research PDF” to refer to the digital version.
-PDF was a derivative of Adobe’s already-successful Postscript programming language for specifying documents, and a proprietary file format until 2008.
+PDF was a derivative of Adobe’s already-successful Postscript language for specifying documents, and a proprietary file format until 2008.
 In an article from 1998, Kasdorf compares the PDF format to SGML (Standard Generalized Markup Language), and argues that both formats should be used on the web: markup languages for screen-based content and page-based languages for printed content (Kasdorf, 1998).
@@
 Didn’t really look very good on the screen.” P7 found the transition to PDF freeing: “I’ve seen figures have become a whole lot more common…the PDFs are now in color so you have a whole bunch of other choices.”
-One such choice is the high resolution figure, which exploits the scalable nature of vector graphics and the zoom feature of PDF viewers to put more data in a research PDF without exceeding the page limit.
+One such choice is the high resolution figure, which exploits the scalable nature of vector graphics and the zoom feature of PDF viewers to put more data in a PDF without exceeding the page limit.
 Several examples of this trend from CVPR 2021 are shown in Figure 6.
@@
 Or is it simply because really…the techniques that we’re describing nowadays are so, so much more complicated?
-So you need them to, you know, to, to put the architecture of the whole thing in detail there.
+So you need them to…put the architecture of the whole thing in detail there.
 Otherwise nobody understands it.” P7, on the other hand, ascribes it to digital submissions: “In ’94, you turned in your paper by actually having a piece of large paper that was probably 24” wide by 30” or something…you had columns of text from LaTeX and then your figures had to be glued in place in what’s called ‘screen,’ right?
@@
 The teaser attracts attention, the dense figures provide details for those who zoom in, and the impression of details for those who do not.
-The digital nature of contemporary research PDFs was a sensitive topic in our interviews.
+The digital nature of contemporary research papers was a sensitive topic in our interviews.
 Half of our participants (P1, P2, P3, P4, P5, and P8) expressed attachment to and nostalgia for their printed papers, and were saddened when asked when they stopped reading on paper.
@@
 From this perspective, digitization is an essential component of commodification, as it separates the research from its paper and thus its role as symbolic object.
-While reading digitally is significantly more environmentally sustainable than printing, shipping and photocopying paper (Aydemir and Özsoy, 2020), many of our participants who primarily read on screens miss the era of physical papers.
+While reading digitally is significantly more environmentally sustainable than printing, shipping, and photocopying paper (Aydemir and Özsoy, 2020), many of our participants who primarily read on screens miss the era of physical papers.
 In contrast, P7 prefers PDFs because they allow him to use a screenreader: “I use an app that reads the pdf aloud to me…I can get most of the paper from that, I still have to read equations.” Despite being an advocate for making computer vision papers more accessible for the community, he explains that conferences are reluctant to put accessible equations in their templates because of its impact on file size: “basically the conferences keep using templates that—so there’s a trade off by adding this stuff to your file, your file gets bigger.
@@
 4.3.2. Faster publication through arXiv
-The digital materiality of the research PDF affords an alternative publication process that came up in several of our interviews (P4,P5,P8): the preprint server arXiv.
+The digital materiality of the PDF affords an alternative publication process that came up in several of our interviews (P4, P5, P8): the preprint server arXiv.
 P5 says arXiv is a major source of anxiety for his students: “So my students…had the unpleasant experience of, you know, finishing a paper when they’re just about to submit and they saw a paper on arXiv that did almost the same thing.
@@
 So I feel like, oh my god, nobody’s even gonna read my beautiful prose!…But also I think because the field progresses faster, papers become obsolete much quicker.
-So it might be reasonable not to spend so much effort on a single paper if you know that in a year it will be obsolete.” In other words, author writes differently in order to better fit the faster publication system, spending less time on introductions in order to spend less effort on a paper which might quickly become obsolete.
+So it might be reasonable not to spend so much effort on a single paper if you know that in a year it will be obsolete.” In other words, the author now writes differently in order to better fit the faster publication system, spending less time on introductions in order to spend less effort on a paper which might quickly become obsolete.
 To summarize, the technologies underlying the publishing process in computer vision have changed rapidly over the past three decades.
-Today, proceedings are published online, and most researchers read research PDFs on a screen, rather than research papers in print.
+Today, proceedings are published online, and most researchers read PDFs on a screen, rather than research papers in print.
 The loss of physical research papers affects readers’ attachment to those documents, as virtual papers cannot hold significance as domestic objects in the same way.
@@
 Using a media archaeology approach, we have described the development of three aspects of the design of the contemporary computer vision research paper.
-First, we saw how teaser images, titles with acronyms and videos advertise the contribution of a paper, and how attention from arXiv and social media has become more important.
-Second, we saw how the results table was introduced for measuring the significance of a contribution, and become a ubiquitous part of the computer vision research paper.
+First, we saw how teaser images, titles with acronyms, and videos advertise the contribution of a paper, and how attention from arXiv and social media has become more important.
+Second, we saw how the results table was introduced for measuring the significance of a contribution, and became ubiquitous.
 Finally, we looked at the transition from paper to PDF as a medium for disseminating research contributions and its new affordances for figure color and density.
@@
 While we observe these changes in computer vision, they may soon appear in other disciplines as well.
-When we view online attention as value-producing labor, we see a clear explanation for our visual trends.
-Authors who write more papers which are easy to understand at a glance, easy to promote on social media, more obviously novel and more significant in the eyes of reviewers attract more attention, generating increased benefit from the labor of their peers.
-This benefit comes in a variety of forms including citations, interest from future students and employees and increased progress on specific research problems.
+When we treat online attention from peers as a kind of labor—something which allows authors to generate value from the research activities of their peers—we see a clear explanation for our visual trends.
+Authors who write more papers which are easy to understand at a glance, easy to promote on social media, more obviously novel, and more significant in the eyes of reviewers attract more attention.
+Attention confers a variety of benefits including citations, interest from future mentors, students, employers, and employees, and increased progress on specific research problems.
+These benefits increase the value of the author’s research.
 Authors seeking attention can develop design innovations, like more readable tables, but the attention economy ultimately privileges large industry labs.
@@
 Instead, paying attention to the state of the discipline is essential for doing career-progressing research of their own.
-While staying current has always been a part of academic work, the growth of the discipline puts pressure on it, increasingly treating researchers like a crowd of interchangeable peers whose attention is needed to make new research prestigious.
+While staying current has always been a part of academic work, the growth of the discipline puts pressure on readers, increasingly treating them like a crowd of interchangeable peers whose attention is needed to make new research prestigious.
 This process resembles proletarianization333In Marxist theory, the process by which people are assimilated into the working class, causing alienation.
@@
 These technologies (some of which originate from CVPR, e.g.
-(Li et al., 2023)) show risk of producing online disinformation (Vykopal et al., 2023), displacing artists and writers (Jiang et al., 2023) and creating problems for future curators and historians regarding provenance and authenticity of historical images (Hazan, 2023).
+(Li et al., 2023)) show risk of producing online disinformation (Vykopal et al., 2023), displacing artists and writers (Jiang et al., 2023), and creating problems for future curators and historians regarding provenance and authenticity of historical images (Hazan, 2023).
 For example, as Jiang et al.
@@
 While bans on AI-generated content may slow the growth in the short-term, they do not resolve the fundamental commodification of attention, and may fail as new tools continue to blur the boundary between human and AI-generated writing.
-The main implication is that ad hoc responses to symptoms of the attention economy, like banning discussion on social media or punishing low effort peer reviews will not change the pressures which are at play.
-Instead, computer vision is faced with a wicked problem (Buchanan, 1992), spanning peer review, hiring, promotion and tenure, in addition to academic discourse, which demands a larger-scale rethinking of the inherent tensions between facilitating scholarly communication and evaluating scholarship.
+The main implication is that ad hoc responses to symptoms of the attention economy, like banning discussion on social media or punishing low effort peer reviews, will not change the pressures which are at play.
+Instead, computer vision is faced with a wicked problem (Buchanan, 1992), spanning peer review, hiring, promotion, and tenure, in addition to academic discourse, which demands a larger-scale rethinking of the inherent tensions between facilitating scholarly communication and evaluating scholarship.
 Removing barriers to research communication appears to speed up the pace of research, but also changes disciplinary practices by forcing scholars into an arms race, competing for the attention of their peers.
@@
 There is also a growing need for tools to help scholars collect and curate relevant literature.
-But rather than make this process as fast and frictionless as possible, we encourage work grounded in “slow design” (Odom et al., 2012) and “slow science” (Stengers, 2018) which attempts to recreate the preciousness, uniqueness and care that our participants ascribed to the papers they photocopied by hand.
+But rather than make this process as fast and frictionless as possible, we encourage work grounded in “slow design” (Odom et al., 2012) and “slow science” (Stengers, 2018) which attempts to recreate the preciousness, uniqueness, and care that our participants ascribed to the hardcopy papers they photocopied by hand.
 We also encourage, in line with the view of attention as labor, top conferences to consider compensating peer reviewers for the essential labor they perform.
@@
 If authors are under pressure to publish more and faster, it is easiest to do that by overstating the significance of completed work.
-There are a range of behaviors which contribute to this problem: from neglecting to explore the limitations of a method, to only showing favorable evaluations, to outright results fabrication.
-While we do not claim that authors are engaging in such behaviors, we do observe an tendency to write papers like advertisements and only minimally discuss the downsides of advertised methods.
-These behaviors become problematic when engineers, stakeholders or researchers in other fields who are unfamiliar with the reality of conference publishing may take the claims in papers at face value, and put minimally tested systems into production based on a general trust of computation and data.
+There are a range of behaviors which contribute to this problem: from neglecting to explore the limitations of a method, to only showing favorable evaluations, to outright fabrication of results.
+While we do not claim that authors are engaging in such behaviors, we do observe a tendency to write papers like advertisements and only minimally discuss the downsides of advertised methods.
+These behaviors become problematic when engineers, stakeholders, and researchers in other fields who are unfamiliar with the reality of conference publishing may take the claims in papers at face value, and put minimally tested systems into production based on a general trust of computation and data.
 6. Conclusion
@@
 (Qian et al., 2019; Chang et al., 2023; Head et al., 2022)) can help streamline this process and reduce the burden on individual scholars, the problem will persist unless scholarly communities address the commodification of scholarly attention.
-Generally, we recommend that conference organizers approach this situation as a systemic problem, and investigate measures to avoid competition between authors, slow the submission and peer review process and search for more equitable and efficient ways of distributing attention and prestige.
+Generally, we recommend that conference organizers approach this situation as a systemic problem, and investigate measures to avoid competition between authors, slow the submission and peer review process, and search for more equitable and efficient ways of distributing attention and prestige.
 Acknowledgements.
-This work was supported in part by a NSF graduate research fellowship.
+This work was supported in part by a National Science Foundation Graduate Research Fellowship.
 The views expressed in this paper are those of the authors and do not necessarily reflect the views of the NSF or any other agencies.References